### v2

# RSFT functors for strong cobordisms and applications

Abstract.
We extend the hierarchy functors of [31] for the case of strong symplectic cobordisms, via deformations with Maurer–Cartan elements. In particular, we prove that the concave boundary of a strong cobordism has finite algebraic planar torsion if the convex boundary does, which yields a functorial proof of finite algebraic planar torsion for contact manifolds admitting strong cobordisms to overtwisted contact manifolds. We also show the existence of contact $3$-folds without strong cobordisms to the standard contact $3$-sphere, that are not cofillable. We also include generalizations of the theory, which appeared in an earlier version of [31], relating our notion of algebraic planar torsion to Latschev–Wendl’s notion of algebraic torsion [25], discussing variations from counting holomorphic curves with general constraints and invariants extracted from higher genera holomorphic curves from an algebraic perspective.
1. Introduction
In [31], the authors introduced the formalism of $BL_{\infty}$ algebras, a framework underlying the algebraic aspects of rational symplectic field theory (RSFT). This setup was used to introduce several invariants for contact manifolds, including the notion of planarity (roughly speaking, from rational curves with a point constraint in symplectizations), as well as algebraic planar torsion (roughly speaking, from rational curves without negative punctures in symplectizations), designed to provide a notion of complexity within the category of all contact manifolds, together with Liouville/exact cobordisms between them as morphisms. Several properties and examples were discussed, and in particular, every level of algebraic planar torsion and planarity was shown to be achieved by a contact manifold. In this paper, we consider the $BL_{\infty}$ formalism in the strong symplectic cobordism category and study the functorial properties of algebraic planar torsion and planarity under strong cobordisms. We also discuss several variations, including the higher genera theory, and the theory with general constraints from an algebraic perspective.
1.1. Hierarchy functors for strong cobordisms
In [31], the authors introduced the notion of algebraic planar torsion $\operatorname{APT}$, which is an analog, in the context of RSFT, of Latschev and Wendl’s algebraic torsion in the context of the full SFT. The finiteness of $\operatorname{APT}$ is an obstruction to strong fillings, and $\operatorname{APT}$ of the convex boundary is at least $\operatorname{APT}$ of the concave boundary in an exact cobordism. By [31, Theorem 3.16, 3.17] and [35, Theorem 1], such monotonicity does not hold for strong cobordisms. However, the following theorem implies that the finiteness of $\operatorname{APT}$ is preserved under a strong cobordism. This answers a question of Latschev and Wendl [25, Question 2] affirmatively in the context of RSFT. The original question can be solved similarly.
Theorem 1.1.
If there is a strong cobordism from $Y$ to $Y^{\prime}$ such that $\operatorname{APT}(Y^{\prime})<+\infty$, then $\operatorname{APT}(Y)<+\infty$. In particular, if $Y^{\prime}$ is algebraically overtwisted, i.e. the contact homology of $Y^{\prime}$ vanishes, then $Y$ has finite algebraic planar torsion.
There are many examples of tight contact manifolds that admit strong cobordisms to overtwisted contact manifolds, e.g. any closed contact $3$-manifold with planar torsion [35]. Theorem 1.1 yields new proof that contact manifolds with planar torsion have finite algebraic planar torsion. Strong cobordisms induce $BL_{\infty}$ morphisms from the RSFT of the convex boundary $Y^{\prime}$ to the RSFT deformed by a Maurer–Cartan element of the concave boundary $Y$. The Maurer–Cartan element is defined by counting rigid genus 00 holomorphic caps with multiple negative punctures in the strong cobordism. The deformed functoriality implies that the deformed RSFT of $Y$ has $\operatorname{APT}$ no larger than $\operatorname{APT}(Y^{\prime})$, and then we can deduce that the undeformed $\operatorname{APT}(Y)$ is also finite, but it could be larger than $\operatorname{APT}(Y^{\prime})$. Maurer-Cartan deformations were studied in various symplectic invariants, starting from bounding cochains in Lagrangian Floer theory by Fukaya, Oh, Ohta, and Ono [18], Maurer-Cartan deformations of $BV_{\infty}$ algebras in the full SFT by Cieliebak and Latschev [11], Maurer-Cartan deformations of $IBL_{\infty}$ algebras in the full SFT by Cieliebak, Fukaya and Latschev [10], Maurer-Cartan deformations of $L_{\infty}$ algebras in symplectic cohomology by Borman, Sheridan, and Varolgunes [2].
Another hierarchy functor introduced in [31] is the planarity $\operatorname{P}$, which measures the tightness of, typically, fillable contact manifolds in the exact cobordism category. Similar to $\operatorname{APT}$, the finiteness of $\operatorname{P}_{\Lambda}$ (the planarity using augmentations over the Novikov field $\Lambda$) of a contact manifold $Y$ is an obstruction to strong semi-fillings, i.e. connected strong fillings $W$ of a disjoint union $Y\sqcup Y^{\prime}$. Understanding obstructions to semi-fillings can be useful for obstructions to fillings, see e.g. [15]. We say that $Y$ is strongly co-fillable if there is a strong semi-filling of $Y$.
Theorem 1.2.
If $\operatorname{P}_{\Lambda}(Y)<+\infty$, then $Y$ is not strongly co-fillable.
There are many non-co-fillable contact manifolds, e.g. $(S^{2n-1},\xi_{\mathrm{std}})$ [29], or more generally planar contact manifolds [16], and iterated planar contact manifolds [1]. All of them have finite planarity [31]. Indeed, the proof of Theorem 1.2 is essentially an RSFT packaging of Eliashberg’s filling by holomorphic disks technique, which forms part of the proof of the celebrated Eliashberg–Floer–McDuff theorem [29].
One would expect that the finiteness of planarity is also preserved under strong cobordisms. The functorial properties of Maurer–Cartan deformed structures in the definition of planarity do hold, hence finiteness of planarity is preserved under strong cobordisms if we use any fixed augmentation of the concave boundary. However, it is still possible that the upper bound blows up when we enumerate all possible augmentations. Hence we can only prove the functoriality of the finiteness of planarity if the convex boundary’s finite planarity is contributed by curves that do not depend on augmentations. This is, in fact, the situation for all examples of finite planarity considered in [31].
Theorem 1.3.
Let $Y^{\prime}$ be a contact manifold with finite planarity, where the contributing curves do not depend on augmentations in the sense of Proposition 4.5 (holds for all explicit examples in [31]), and let $X$ be a strong cobordism from $Y$ to $Y^{\prime}$. Then $\operatorname{P}(Y)\leq\operatorname{P}_{\Lambda}(Y)<+\infty$.
As a corollary, we have the following explicit examples, taken from [31].
Corollary 1.4.
Let $s$ be a holomorphic section of ${\mathcal{O}}(d)$ for $d\geq 1$ over $\mathbb{C}\mathbb{P}^{n}$. Let $Y$ be the contact boundary of the affine variety $\mathbb{C}\mathbb{P}^{n}\backslash(s^{-1}(0)\cup H)$, where $H$ is a hyperplane. Then $\operatorname{P}(Y)\leq\operatorname{P}_{\Lambda}(Y)\leq d+1$.
Remark 1.5.
Let $D_{1},D_{2}$ be two ample divisors of $\mathbb{C}\mathbb{P}^{n}$. Naively, one may expect that the planarity of $\partial(\mathbb{C}\mathbb{P}^{n}\backslash D_{1})$ is not larger than that of $\partial(\mathbb{C}\mathbb{P}^{n}\backslash(D_{1}\cup D_{2}))$, as many computations in [31] indicated. Combining Corollary 1.4 and [31, Theorem 7.13], we know that such expectation does not hold in general, e.g. for $D_{1}$ a smooth hypersurface of sufficiently high degree and $D_{2}=H$ in $\mathbb{C}\mathbb{P}^{n}$ for $n\geq 2$.
Lastly, we recall that, given an open book supporting a contact structure on $M$, Bourgeois [4] constructed a contact structure on $M\times\mathbb{T}^{2}$. Several applications of this construction were obtained in [7, 8]. For these manifolds, we can estimate their complexity in general.
Corollary 1.6.
The planarity (over ${\mathbb{Q}}$ or $\Lambda$) of any Bourgeois contact manifold is at most $2$.
We also make the following observation that there exist contact $3$-manifolds without strong cobordism to $(S^{3},\xi_{\mathrm{std}})$ that are not cofillable, which settles a question of Wendl [35, Question 2] in the negative.
Theorem 1.7.
For $g\geq 1$ and $d>2g-2$, let $Y_{d,g}$ be the pre-quantization bundle, equipped with the Boothby-Wang contact structure, over a surface $\Sigma_{g}$ of genus $g$ with first Chern class $-d$. Then $Y_{d,g}$ is not strongly co-fillable and there is no strong cobordism with concave boundary $Y_{d,g}$ and convex boundary $(S^{3},\xi_{\mathrm{std}})$.
1.2. Higher genera and other generalizations
In [20], Gay asked the question that, under the equivalence relation defined by the existence of strong cobordisms in both directions, whether there is an equivalence class of contact $3$-manifolds beyond []delimited-[][$\emptyset$][ ∅ ] and $[(S^{3},\xi_{\mathrm{ot}})]$, where $\xi_{\mathrm{ot}}$ is an overtwisted contact structure. This question is equivalent to whether there is a contact $3$-manifold that is not strongly fillable and has no strong cobordism to an overtwisted manifold. The same question was also asked by Wendl [35, Question 1]. The answer is affirmative by contact manifolds with non-vanishing Heegaard-Floer/monopole-Floer/ECH contact invariant that are not strongly fillable [27]. This follows from the fact that the vanishing of contact invariant is preserved under strong cobordisms by [12] (in monopole-Floer) and [22] (in ECH).
If we approach this question (and its higher dimensional analog) from the SFT perspective, the functoriality in Theorem 1.1 implies that a non-strongly-fillable contact manifold with infinite $\operatorname{APT}$ can also serve as a solution. Therefore we need obstructions to strong fillings beyond the finiteness of $\operatorname{APT}$. The obvious direction is upgrading the invariants from RSFT to those from the full SFT, i.e. considering curves with all genera. The counterpart of $\operatorname{APT}$ in the full SFT is the algebraic torsion $\mathrm{AT}$ in [25]. Although both torsions share the same class of examples of $3$-manifolds with finite torsions, they do not have any obvious direct relation, at least algebraically. In fact, they are related by a grid of torsions filtered by both the number of punctures and genera. However, finding examples realizing those torsions is a challenging task, as we need to find certain higher genus curves in the contact manifold. In §5, which was contained in an earlier version of [31], we discuss the higher genera analogs of our hierarchy functors, and the relation between two notions of torsions (see also [24] for a reformulation of our invariants in the language of the original description of SFT by Eliashberg, Givental and Hofer [13]). We will also discuss several variations, in the spirit of Siegel’s higher capacities [34], on invariants coming from counting rational curves with multiple point constraints, as well as more general constraints for our future applications.
Acknowledgements. A. Moreno is supported by the National Science Foundation under Grant No. DMS-1926686, by the Sonderforschungsbereich TRR 191 Symplectic Structures in Geometry, Algebra and Dynamics, funded by the DFG (Projektnummer 281071066 – TRR 191), and also by the DFG under Germany’s Excellence Strategy EXC 2181/1 - 390900948 (the Heidelberg STRUCTURES Excellence Cluster). Z. Zhou is supported by the National Key R&D Program of China under Grant No. 2023YFA1010500, the National Natural Science Foundation of China under Grant No. 12288201 and 12231010.
2. Algebraic formalism
2.1. $BL_{\infty}$ algebras
We start by recalling the main algebraic notions from [31], following the same (but streamlined) exposition. Let $V$ be a ${\mathbb{Z}}/2$-graded vector space over a field ${\bm{k}}$ of characteristic zero. Consider the ${\mathbb{Z}}/2$-graded symmetric algebra $SV:=\bigoplus_{k\geq 0}S^{k}V$ and the non-unital symmetric algebra $\overline{S}V=\bigoplus_{k\geq 1}S^{k}V$, where $S^{k}V=\otimes^{k}V/\Sigma_{k}$ in the graded sense, i.e.
$ab=(-1)^{|a||b|}ba$
for homogeneous elements $a,b$ in $SV,\overline{S}V$.
Let $EV=\overline{S}SV$. We use $\odot$ for the product on the outside symmetric product $\overline{S}$ and $\ast$ (often omitted) for the product on the inside symmetric product $S$. Given the data of a linear operator $p^{k,l}:S^{k}V\to S^{l}V$ for $k\geq 1,l\geq 0$, we define a map $\widehat{p}^{k,l}:S^{k}SV\to SV$, by the following properties.
  1. (1)
    $\widehat{p}^{k,l}|_{\odot^{k}V\subset S^{k}SV}$ is defined by $p^{k,l}$.
  2. (2)
    If $w_{i}\in{\bm{k}}$, then $\widehat{p}^{k,l}(w_{1}\odot\ldots\odot w_{k})=0$.
  3. (3)
    $\widehat{p}^{k,l}$ satisfies the Leibniz rule in each argument, i.e. we have
    (2.1) $\widehat{p}^{k,l}(w_{1}\odot\ldots\odot w_{k})=\sum_{j=1}^{m}(-1)^{\square}v_{1}\ldots v_{j-1}\widehat{p}^{k,l}(w_{1}\odot\ldots\odot v_{j}\odot\ldots\odot w_{k})v_{j+1}\ldots v_{m}.$
    Here $w_{i}=v_{1}\ldots v_{m}$ and
    (2.2) $\square=\sum_{s=1}^{i-1}|w_{s}|\cdot\sum_{s=1}^{j-1}|v_{s}|+\sum_{s=1}^{j-1}|v_{s}||p^{k,l}|+\sum_{s=i+1}^{n}|w_{s}|\cdot\sum_{s=j+1}^{m}|v_{s}|.$
Explicitly, $\widehat{p}^{k,l}$ is defined by
(2.3) $w_{1}\odot\ldots\odot w_{k}\mapsto\sum_{\begin{subarray}{c}(i_{1},\ldots,i_{k})\\ 1\leq i_{j}\leq n_{j}\end{subarray}}(-1)^{\bigcirc}p^{k,l}(v^{1}_{i_{1}}\odot\ldots\odot v^{k}_{i_{k}})\check{w}_{1}\ldots\check{w}_{k},$
where $w_{j}=v^{j}_{1}\ldots v^{j}_{n_{j}}$, $\check{w}_{j}=v^{j}_{1}\ldots\check{v}^{j}_{i_{j}}\ldots v^{j}_{n_{j}}$ and $w_{1}\ldots w_{k}=(-1)^{\bigcirc}v^{1}_{i_{1}}\ldots v^{k}_{i_{k}}\check{w}_{1}\ldots\check{w}_{k}$. Then we define $\widehat{p}^{k}:S^{k}SV\to SV$ by $\bigoplus_{l\geq 0}\widehat{p}^{k,l}$. Then we can define $\widehat{p}:EV\to EV$ by
(2.4) $w_{1}\odot\ldots\odot w_{n}\mapsto\sum_{k=1}^{n}\sum_{\sigma\in Sh(k,n-k)}(-1)^{\diamond}\widehat{p}^{k}(w_{\sigma(1)}\odot\ldots\odot w_{\sigma(k)})\odot w_{\sigma(k+1)}\odot\ldots\odot w_{\sigma(n)},$
where $Sh(k,n-k)$ is the subset of permutations $\sigma$ such that $\sigma(1)<\ldots<\sigma(k)$ and $\sigma(k+1)<\ldots<\sigma(n)$ and
$\diamond=\sum_{\begin{subarray}{c}1\leq i<j\leq n\\ \sigma(i)>\sigma(j)\end{subarray}}|w_{i}||w_{j}|.$
Definition 2.1 ([31, Definition 2.3]).
$(V,\{p^{k,l}\})$ is a $BL_{\infty}$ algebra if $\widehat{p}\circ\widehat{p}=0$ and $|\widehat{p}|=1$.
The definition above implicitly contains the requirement that $\widehat{p}$ is well-defined. To ensure such, for example, we can impose that for any $v_{1},\ldots,v_{k}\in V$, there are at most finitely many $l$ such that $p^{k,l}(v_{1}\odot\ldots\odot v_{k})\neq 0$.
2.2. The rules for tree calculus
A useful way to explain the combinatorics of operations is the following description using graphs, which appeared in [34, §3.4.2]. The main advantage of such graphical language is freeing us from the book-keeping of signs and explicit components of compositions, e.g. in (2.1), which are governed by graphs.
Let $w\in S^{k}V$, we can represent $w$ by an element $\overline{w}$ in $\otimes^{k}V$, that is $\overline{w}=\sum_{i=1}^{N}c_{i}v^{i}_{1}\otimes\ldots\otimes v^{i}_{k}$ for $c_{i}\in{\bm{k}}$ and $v_{*}^{*}\in V$, such that $\pi(\overline{w})=w$ for $\pi:\otimes^{k}V\to S^{k}V$. We represent it by a rooted tree with $k$ leaves (represented by $\bullet$) labeled by $\overline{w}$. The leaves are ordered from left to right to indicate the $k$ copies of $V$ in $\otimes^{k}V$. When $\overline{\omega}=v_{1}\otimes\ldots\otimes v_{k}$, we may label the leaves by $v_{1},\ldots,v_{k}$ to mean the same thing. We can view a general labeled tree as a formal linear combination of such trees with leaves labeled.
$\overline{w}\in\otimes^{3}V$$v_{1}$$v_{2}$$v_{3}$
Now let $s\in S^{k}SV$, we can represent $s$ by $\overline{s}\in\boxtimes^{k}TV$, where $TV=\oplus_{k\in{\mathbb{N}}}(\otimes^{k}V)$. Here we use $\boxtimes$ to differentiate it from the inner tensor $\otimes$. We write
$\overline{s}=\sum_{i=1}^{N}c_{i}\overline{w}^{i}_{1}\boxtimes\ldots\boxtimes\overline{w}^{i}_{k},\quad c_{i}\in{\bm{k}},\overline{w}_{*}^{*}\in\otimes^{m^{*}_{*}}V.$
We represent $\overline{w}^{i}_{1}\boxtimes\ldots\boxtimes\overline{w}^{i}_{k}$ by an ordered forest of labeled trees as follows. Then $\overline{s}$ is a formal linear combination of such forests.
$\overline{w}_{1}\in\otimes^{3}V$$\overline{w}_{2}\in\otimes^{3}V$$\overline{w}_{3}\in\otimes^{2}V$
Figure 1. forest of labeled trees
We represent the operation $p^{k,l}:S^{k}V\to S^{l}V$ by a graph with $k+l+1$ vertices, $k$ top input vertices, $l$ bottom output vertices and one middle vertex labeled by $p^{k,l}$ representing the operation type.
$p^{2,3}$
So far the discussion is completely formal without any actual content, the real content is in the following interpretation of a glued graph, whose definition will be clear from one example.
$v_{1}$$v_{2}$$v_{3}$$v_{4}$$v_{5}$$v_{6}$$v_{7}$$v_{8}$
Figure 2. Gluing forests $\Leftrightarrow$ applying operations
The above glued graph represents a forest: we first fix a representative $\overline{p}$ of $p^{2,3}(v_{3}v_{4})$ in $\otimes^{3}V$. The glued forest in Figure 1 represents ±$\pm(v_{1}\otimes v_{2}\otimes\overline{p}\otimes v_{5}\otimes v_{6})\boxtimes(v_{7}\otimes v_{8})$. In the gluing, we do not create cycles in the glued graph, each dashed line represents the identity map, and each connected component represents a tree in the output. Drawing the input element as a forest of ordered trees with ordered leaves corresponding to that we are choosing representatives from the tensor product, not the symmetric product. Finally, when we draw the glued graph on a plane as above, i.e. choosing an order of the trees and leaves (hence edges will cross over each other), will determine a representative in the tensor product, but we view different orders as equivalent up to the obvious sign change. For example, the following is an equivalent gluing with Figure 1 but with an extra sign when viewing it in the tensor product.
$v_{1}$$v_{2}$$v_{3}$$v_{4}$$v_{5}$$v_{6}$$v_{7}$$v_{8}$
The sign is determined similarly to (2.2), in this case of Figure 2, the sign is $(-1)^{(|v_{1}|+|v_{2}|)|p^{2,3}|}$. In a formal description, we apply order changes to the input forest (edges can cross), then we glue $p^{k,l}$, such that there is no new edge crossing, finally, we change the output order back to the chosen one, the final sign is given by the product of the sign changes of the two order changes and the sign of the composition using the Koszul-Quillen convention. Writing the forest using a glued graph as in Figure 2 contains slightly more refined information than just labeling the forest as in Figure 1, namely we keep track of which leaves are from $p^{k,l}$ in a representative. The following observation is tautological.
Proposition 2.2.
The output of a glued forest is well-defined in $EV$.
To enumerate through all admissible gluings, each output leaf and tree are considered as different. However, we do not differentiate the input leaves of $p^{k,l}$. Therefore when we glue a $p^{k,l}$ component, we pick $k$ trees (this is $Sh(k,n-k)$ in (2.4) ) from the forest and then one leaf from each chosen tree (that is $1\leq i_{j}\leq n_{j}$ in (2.3)) to glue to $p^{k,l}$. For example, in the situation of Figure 2, we have $3*3+3*2+3*2=21$ direct ways to glue $p^{2,3}$. The ambiguity from choosing a representative of the input is then eliminated from summing over all possible gluing by the following tautological observation.
Proposition 2.3.
When summed over all possible gluing of one $p^{k,l}$, the output is independent of the choice of representatives of the input forest.
Combining the above two propositions, we see that gluing forests corresponds to operations on $EV$. Indeed, the language of trees and forests is just packaging the signs and components in (2.1), (2.3) and (2.4) by providing a geometric intuition. The translation into forests makes it easier to understand algebraic relations, for example, many relations come from interpreting the same glued forests in two different ways. The following is the dictionary of the algebraic formulae in §2.1 in terms of forests.
$\widehat{p}^{k,l}$ on $\odot^{k}V$ (unique) gluing of $p^{k,l}$ to a forest of $k$ trees of single leaf
$\widehat{p}^{k,l}$ in (2.3) sum of gluing of $p^{k,l}$ to a forest of $k$ trees to get a tree
$\widehat{p}^{k}$ sum of gluing of $p^{k,*}$ to a forest of $k$ trees to get a tree
$\widehat{p}$ in (2.4) sum of gluing of $p^{*,*}$ to a forest to get a forest
$\widehat{p}^{k,l}(w_{1}\odot\ldots\odot w_{k})=0$ if $w_{i}\in{\bm{k}}$ no way to glue $p^{k,l}$ to a forest with a tree without leaf to get a tree
In the following, we will use the graphical description for morphisms, augmentations, and Maurer-Cartan elements without giving the explicit formulae like (2.1), (2.4). The explicit formulae, except for the Maurer-Cartan elements, can be found in [31, §2].
2.3. Morphisms
We now define morphisms between $BL_{\infty}$ algebras. Consider a family of operators $\{\phi^{k,l}:S^{k}V\to S^{l}V^{\prime}\}_{k\geq 1,l\geq 0}$, we can construct a map $\widehat{\phi}:EV\to EV^{\prime}$ from the following tree description. To represent $\phi^{k,l}$, we use a graph similar to the one representing $p^{k,l}$ but replace by to indicate that they are maps of different roles. To represent a component configuration in the definition of $\widehat{\phi}$ on $S^{i_{1}}V\odot\ldots\odot S^{i_{n}}V$, we glue a family of graphs representing $\phi^{k,l}$ such that input vertices and the output vertices of the top forest are completely paired and glued and the resulted graph has no cycles. Then $\widehat{\phi}$ is the sum of all possible configurations. Unlike the definition of $\widehat{p}$ that we need to glue exactly one $p^{k,l}$ graph, it is possible that we do not glue in any $\phi^{k,l}$ graphs. This is the case when the input in $\odot^{m}{\bm{k}}$ and $\widehat{\phi}$ is identity in such case, i.e. $\widehat{\phi}(1\odot\ldots\odot 1)=1\odot\ldots\odot 1$.
Figure 3. A component of $\widehat{\phi}$ from $S^{3}V\odot S^{3}V\odot S^{2}V$ to $S^{6}V^{\prime}$
Definition 2.4 ([31, Definition 2.9]).
$\widehat{\phi}$ is a $BL_{\infty}$ morphism from $(V,\widehat{p})$ to $(V^{\prime},\widehat{p}^{\prime})$ if $\widehat{\phi}\circ\widehat{p}=\widehat{p}^{\prime}\circ\widehat{\phi}$ and $|\widehat{\phi}|=0$.
The definition above implicitly assumes that $\widehat{\phi}$ is well-defined. For example, when $\phi$ arises from counting holomorphic curves in an exact cobordism, we have for any $v_{1}\ldots v_{k}\in S^{k}V$ there are at most finitely many $l$, such that $\phi^{k,l}(v_{1}\ldots v_{k})\neq 0$. Hence $\widehat{\phi}$ is well-defined.
The $(\psi\circ\phi)^{k,l}$ component of the composition of two $BL_{\infty}$ morphisms $\phi,\psi$ is represented by connected graphs without cycles glued from one level from $\phi$ and one level from $\psi$. It is clear from the tree description that $\widehat{\psi\circ\phi}=\widehat{\psi}\circ\widehat{\phi}$ and the composition is associative, see [31, §2.4].
2.4. Augmentations
The trivial vector space $V=\{0\}$ has a unique trivial $BL_{\infty}$ algebra structure by $p^{k,l}=0$. We use $\mathbf{0}$ to denote this trivial $BL_{\infty}$ algebra.
Definition 2.5 ([31, Definition 2.11]).
A $BL_{\infty}$ augmentation is a $BL_{\infty}$ morphism $\epsilon:V\to\mathbf{0}$, i.e. a family of operators $\epsilon^{k}:S^{k}V\to{\bm{k}}$ satisfying Definition 2.4.
For a $BL_{\infty}$ algebra $V$, we let $E^{k}V=\overline{B}^{k}SV:=\bigoplus_{j=1}^{k}S^{j}SV$, which is a filtration (sentence length filtration) on $EV$ compatible with the differential $\widehat{p}$. We have $E\mathbf{0}={\bm{k}}\oplus S^{2}{\bm{k}}\oplus+\ldots$ with $\widehat{p}=0$, and $H_{*}(E^{k}\mathbf{0})=E^{k}\mathbf{0}$ for all $k\geq 1$. We define $1_{\mathbf{0}}$ be the generator in $E^{1}\mathbf{0}$, then $1_{\mathbf{0}}\neq 0\in H_{*}(E^{k}\mathbf{0})$ for all $k\geq 1$. Then we define $1_{V}\in H_{*}(E^{k}V)$ to be the image of $1_{\mathbf{0}}$ under the chain map $E^{k}\mathbf{0}\to E^{k}V$ induced by the trivial $BL_{\infty}$ morphism $\mathbf{0}\to V$.
Definition 2.6 ([31, Definition 2.13]).
We define the torsion of a $BL_{\infty}$ algebra $V$ to be
$\operatorname{T}(V):=\min\{k-1|1_{V}=0\in H_{*}(E^{k}V),k\geq 1\}.$
Here the minimum of an empty set is defined to be $\infty$.
By definition, we have that $\operatorname{T}(V)=0$ iff $1_{V}\in H^{*}(SV,\widehat{p}^{1})$ is zero. Since $H^{*}(SV,\widehat{p}^{1})$ is an algebra with $1_{V}$ a unit, in this case we have $H^{*}(SV,\widehat{p}^{1})=0$. In general, a finite torsion is an obstruction to augmentations, and the hierarchy of torsion is functorial w.r.t. $BL_{\infty}$ morphisms.
Proposition 2.7 ([31, §2.5]).
If there is a $BL_{\infty}$ morphism from $V$ to $V^{\prime}$, then $\operatorname{T}(V)\geq\operatorname{T}(V^{\prime})$. In particular, if $\operatorname{T}(V)<\infty$, there is no augmentation (as $\operatorname{T}(\mathbf{0})=+\infty$).
Given a $BL_{\infty}$ augmentation $\epsilon$, we can define a linearized theory via graphs as follows.
$\epsilon^{2}$$\epsilon^{2}$$p^{2,3}$
Figure 4. A component of $p^{4,1}_{\epsilon}$
The linearized structural map $p^{k,l}_{\epsilon}$ is defined by gluing the level containing a $p^{*,*}$ in the definition of $\widehat{p}$ to the bottom level of the definition of $\widehat{\epsilon}$, such that the glued graph is a tree with $k$ inputs and $l$ outputs.
Proposition 2.8 ([31, Proposition 2.16]).
$\{p^{k,l}_{\epsilon}\}$ is a $BL_{\infty}$ algebra structure on $V$ and $p^{k,0}_{\epsilon}=0$ for all $k>0$.
$\{p^{k,l}_{\epsilon}\}$ is $\{p^{k,l}\}$ under a change of coordinates in the $BL_{\infty}$ sense. Namely, we let $F^{1,1}_{\epsilon}=\operatorname{id}_{V}$ and $F^{k,0}_{\epsilon}=\epsilon^{k}$ and all other $F^{k,l}_{\epsilon}=0$. The same recipe of constructing $\widehat{\phi}$ from $\phi^{k,l}$ applies to give a $\widehat{F}_{\epsilon}$ on $EV$. If $\widehat{F}_{-\epsilon}$ denotes the map on $EV$ defined by $F^{1,1}_{-\epsilon}=\operatorname{id}_{V}$ and $F^{k,0}_{-\epsilon}=-\epsilon^{k,0}$ and all other $F^{k,l}_{-\epsilon}=0$, then $\widehat{F}_{-\epsilon}$ is the inverse of $\widehat{F}_{\epsilon}$. Then we have $\widehat{p}_{\epsilon}:=\widehat{F}_{\epsilon}\circ\widehat{p}\circ\widehat{F}_{-\epsilon}:EV\to EV$, see discussions before [31, Proposition 2.16].
2.5. Pointed maps
Let $p_{\bullet}^{k,l}:S^{k}V\to S^{l}V,k\geq 1,l\geq 0$ be a family of linear maps, we can define $\widehat{p}_{\bullet}$ just like $\widehat{p}$ by gluing trees with the modification that the grading $|p_{\bullet}^{k,l}|$ is not necessarily $21\in{\mathbb{Z}}/$ 2.
Definition 2.9 ([31, Definition 2.17]).
We say $\{p_{\bullet}^{k,l}\}$ is a pointed map, iff $\widehat{p}_{\bullet}\circ\widehat{p}=(-1)^{|\widehat{p}_{\bullet}|}\widehat{p}\circ\widehat{p}_{\bullet}$.
In the context of SFT, $p_{\bullet}$ comes from counting holomorphic curves in the symplectization with a marked point mapped to a closed submanifold/closed chain. The degree of $p_{\bullet}$ is the dimension of constraint. In applications, we typically consider a point constraint and $|p_{\bullet}|=0$. If we have an augmentation $\epsilon$ to $p$, we then define
$\widehat{p}_{\bullet,\epsilon}:=\widehat{F}_{\epsilon}\circ\widehat{p}_{\bullet}\circ\widehat{F}_{-\epsilon}.$
This map is determined by the $p^{k,l}_{\bullet,\epsilon}$, which is defined similarly to $p^{k,l}_{\epsilon}$. We also have $\widehat{p}_{\epsilon}\circ\widehat{p}_{\bullet,\epsilon}=\widehat{p}_{\bullet,\epsilon}\circ\widehat{p}_{\epsilon}$. However, it is not necessarily true that $p^{k,0}_{\bullet,\epsilon}=0$. We let $\ell_{\bullet,\epsilon}^{k,0}:=p^{k,0}_{\bullet,\epsilon}$. Then ^$\widehat{\ell}_{\bullet,\epsilon}:=\sum_{k\geq 0}\ell_{\bullet,\epsilon}^{k,0}$ defines a chain morphism $(\overline{S}V,\widehat{\ell}_{\epsilon})\to{\bm{k}}$, where ^$\widehat{\ell}_{\epsilon}$ is the restriction of $\widehat{p}_{\epsilon}$ on $\overline{S}V\cong\bigoplus_{i=1}^{\infty}\odot^{i}V\subset EV$ then composed with the projection to $\bigoplus_{i=1}^{\infty}\odot^{i}V$. In other words, ^$\widehat{\ell}_{\epsilon}$ is determined by $\ell^{k}_{\epsilon}:=p^{k,1}_{\epsilon}$ by a similar recipe of $\widehat{p}$, i.e. the differential on the reduced bar complex of the $L_{\infty}$ algebra $\ell^{k}_{\epsilon}$ on $V[-1]$, see [31, §2.1].
Definition 2.10 ([31, Definition 2.18]).
Given a $BL_{\infty}$ augmentation and a pointed morphism $p_{\bullet}$, the $(\epsilon,p_{\bullet})$-order of $V$ is defined to be
$O(V,\epsilon,p_{\bullet}):=\min\left\{k\left|1\in\operatorname{im}\widehat{\ell}_{\bullet,\epsilon}|_{H_{*}(\overline{S}^{k}V,\widehat{\ell}_{\epsilon})}\right.\right\},$
where the minimum of an empty set is defined to be $\infty$.
To discuss the functoriality of the order above, we need to introduce the notion of morphisms between pointed maps. Given a $BL_{\infty}$ morphism $\phi:(V,p)\to(V^{\prime},q)$ and a family of morphisms $\phi^{k,l}_{\bullet}:S^{k}V\to S^{l}V^{\prime}$, we can define $\widehat{\phi}_{\bullet}:EV\to EV^{\prime}$ by the same rule of $\widehat{\phi}$ with exactly one $\phi^{k,l}_{\bullet}$ component and all the others are $\phi^{k,l}$ components.
Definition 2.11 ([31, Definition 2.19]).
Assume $p_{\bullet},q_{\bullet}$ are two pointed morphisms of $(V,p),(V^{\prime},q)$ respectively, of the same degree. We say $p_{\bullet},q_{\bullet},\phi$ are compatible, if there is a family of $\phi^{k,l}_{\bullet}$ such that $\widehat{q}_{\bullet}\circ\widehat{\phi}-(-1)^{|\widehat{q}_{\bullet}|}\widehat{\phi}\circ\widehat{p}_{\bullet}=\widehat{q}\circ\widehat{\phi}_{\bullet}-(-1)^{|\widehat{\phi}_{\bullet}|}\widehat{\phi}_{\bullet}\circ\widehat{p}$ and $|\widehat{\phi}_{\bullet}|=|\widehat{p}_{\bullet}|+1$.
The following proposition is the functoriality of the $(\epsilon,p_{\bullet})$-order.
Proposition 2.12 ([31, Proposition 2.20]).
Assume $\phi$ is a $BL_{\infty}$-morphism from $(V,p)$ to $(V^{\prime},q)$ with pointed maps $p_{\bullet},q_{\bullet}$ of degree 00 respectively, such that $p_{\bullet},q_{\bullet},\phi$ are compatible. Then for any $BL_{\infty}$ augmentation $\epsilon$ of $V^{\prime}$, we have $O(V,\epsilon\circ\phi,p_{\bullet})\geq O(V^{\prime},\epsilon,q_{\bullet})$.
Remark 2.13.
In view of the discussion above, from the perspective of $(\epsilon,p_{\bullet})$-order, we should view $(p,p_{\bullet})$ as an object, i.e. $p_{\bullet}$ is an additional structure on the $BL_{\infty}$ algebra and $(\phi,\phi_{\bullet})$ as a morphism. This is the reason why we call $p_{\bullet}$ a pointed map instead of morphism, besides the different combinatorial nature compared to a $BL_{\infty}$ morphism.
2.6. Completions and twisted coefficients
Definition 2.14.
We define the Novikov ring $\Lambda_{0}$ over a field ${\bm{k}}$ as
$\Lambda_{0}=\left\{\sum_{i=1}^{\infty}a_{i}T^{\lambda_{i}}\left|a_{i}\in{\bm{k}},\lambda_{i}\geq 0,\displaystyle\lim_{i\rightarrow\infty}\lambda_{i}=+\infty\right.\right\},$
and the Novikov field $\Lambda$ is defined to be
$\Lambda=\left\{\sum_{i=1}^{\infty}a_{i}T^{\lambda_{i}}\left|a_{i}\in{\bm{k}},\displaystyle\lim_{i\to\infty}\lambda_{i}=+\infty\right.\right\}.$
Let $\omega:G\to\mathbb{R}$ be a group homomorphism. The Novikov completion $\overline{{\bm{k}}[G]}$ of the group ring ${\bm{k}}[G]$ is
$\overline{{\bm{k}}[G]}=\left\{\sum_{i=1}^{\infty}a_{i}T^{g_{i}}\left|a_{i}\in{\bm{k}},\displaystyle\lim_{i\to\infty}\omega(g_{i})=+\infty\right.\right\}.$
Those coefficient rings are equipped with a decreasing filtration and are complete w.r.t. the filtration. Take $\Lambda$ as an example: we define $\Lambda_{\rho}$ to be those elements with $\lambda_{i}\geq\rho$, then we have $\Lambda_{\rho_{1}}\subset\Lambda_{\rho_{2}}$ if $\rho_{1}\geq\rho_{2}$ and $\Lambda_{\rho_{1}}\cdot\Lambda_{\rho_{2}}\subset\Lambda_{\rho_{1}+\rho_{2}}$.
Now let $V$ be a ${\bm{k}}$-vector space with filtration degree 00, then $V\otimes_{{\bm{k}}}\Lambda_{0},V\otimes_{{\bm{k}}}\Lambda$ have induced decreasing filtration, and so do the base changes for $SV,EV$. We use $\overline{V\otimes_{{\bm{k}}}\Lambda_{0}},\overline{SV\otimes_{{\bm{k}}}\Lambda_{0}}$, etc. to denote the corresponding completions. For example
$\overline{V\otimes_{{\bm{k}}}\Lambda_{0}}=\left\{\sum_{i=1}^{\infty}v_{i}T^{\lambda_{i}}\left|v_{i}\in V,\lambda_{i}\geq 0,\lim_{i\to\infty}\lambda_{i}=+\infty\right.\right\}$
Definition 2.15.
A filtered completed $BL_{\infty}$ algebra structure on $V$ over $R=\Lambda_{0},\Lambda$ consists of maps $p^{k,l}:\overline{S^{k}V\otimes_{{\bm{k}}}R}\to\overline{S^{l}V\otimes_{{\bm{k}}}R}$ for $k\geq 1,l\geq 0$, such that the assembled map $\widehat{p}$ is well-defined on $\overline{EV\otimes_{{\bm{k}}}R}$ of degree $1$, squares to zero and preserves the filtration. The filtered completed version for morphisms, augmentations, pointed maps and compatible pointed morphisms are similar.
2.7. Maurer-Cartan elements
Although one can define Maurer-Cartan elements for uncompleted algebras, they typically arise in the completed theory in the context of SFT, where they are defined by counting holomorphic curves without positive punctures in a strong cobordism. That they are defined in the completed theory is a consequence of the SFT compactness.
Definition 2.16.
A Maurer-Cartan element of a filtered completed $BL_{\infty}$ algebra is an element $\mathfrak{mc}$ of degree 00 with a positive filtration degree in $\overline{SV\otimes_{{\bm{k}}}R}$, such that $\widehat{p}(e^{\mathfrak{mc}}-1)=0$, where $e^{\mathfrak{mc}}=\sum_{i=0}^{\infty}\frac{\odot^{i}\mathfrak{mc}}{i!}\in\overline{SSV\otimes_{{\bm{k}}}R}=\overline{R}\oplus\overline{EV\otimes_{{\bm{k}}}R}.$ It is clear that $e^{\mathfrak{mc}}-1$ is the projection of $e^{\mathfrak{mc}}$ to $\overline{EV\otimes_{{\bm{k}}}R}$.
Given $a\in\overline{SV\otimes_{{\bm{k}}}R}$ of degree 00 with a positive filtration degree, we define
$\exp_{a}:\overline{EV\otimes_{{\bm{k}}}R}\to\overline{EV\otimes_{{\bm{k}}}R},x\mapsto x\odot e^{a}$
which preserves the filtration.
[…]
We can represent an element $x$ of degree 00 of $\overline{SV\otimes_{{\bm{k}}}R}$ by a sum of labeled trees, i.e. $x=\sum T_{k}$ for $T_{k}\in\overline{S^{k}V\otimes_{{\bm{k}}}R}$
Here we use a notation similar to a $BL_{\infty}$ morphism $\phi$, i.e. trees with representing the root, instead of trees with undotted roots in the description of elements in $SV$, since Maurer-Cartan elements behave like a “morphism” from the trivial $BL_{\infty}$ algebra. We can represent the constant term of $\mathfrak{mc}$ by a tree without leaves, i.e. a single root. With such notation, $e^{x}$ in $\overline{EV\otimes_{{\bm{k}}}R}$ is represented by linear combinations of forests generated by the labeled trees. As each component of $\mathfrak{mc}$ has degree 00, forests up to switching order are equivalent.
Here the coefficient of a forest $\{\underbrace{T_{a_{1}},\ldots,T_{a_{1}}}_{i_{1}},\ldots,\underbrace{T_{a_{m}},\ldots,T_{a_{m}}}_{i_{m}}\}$ is given by the “isotropy order” $\frac{1}{(i_{1})!\ldots(i_{m})!}$. Given a $BL_{\infty}$ morphism $\phi$, then we can define $\phi(x)$ by gluing $\widehat{\phi}$ to the forest above (with the above coefficient) to get a tree, i.e.
Lemma 2.17.
For an element $x$ of degree 00 with a positive filtration degree in $\overline{SV\otimes_{{\bm{k}}}R}$ and $\phi$ a filtered completed $BL_{\infty}$ morphism, we have
(2.5) $e^{\phi(x)}=\widehat{\phi}(e^{x}).$
Proof.
It is clear that both sides of (2.5) count the same type of forests, it suffices to verify the coefficients match. We can look at the coefficient from $T_{k}$. Assume the glued forest has $m_{i}$ identically labeled trees each involving $n_{i}$ trees of $T_{k}$ for $1\leq i\leq p$. Then from the $e^{\phi(x)}$ side, the coefficient from $T_{k}$ is
$\prod_{i=1}^{p}\frac{1}{m_{i}!}\left(\frac{1}{n_{i}!}\right)^{m_{i}}.$
From the $\widehat{\phi}(e^{x})$ side, the coefficient from $T_{k}$ is
$\frac{C}{\displaystyle\left(\sum_{i=1}^{p}m_{i}n_{i}\right)!}$
where $C$ is the number of ways to write $(1,\ldots,\sum_{i=1}^{p}m_{i}n_{i})$ as $p$ ordered subsets with $m_{i}n_{i}$ elements for $1\leq i\leq p$, such that each subset is a union of $m_{i}$ subsets of $n_{i}$ elements. This number is the number of ways of grouping $T_{k}$ in $e^{x}$ so that the glued forest is the one above. Since $C$ is precisely
$\frac{\displaystyle\left(\sum_{i=1}^{p}m_{i}n_{i}\right)!}{\displaystyle\prod_{i=1}^{p}m_{i}!(n_{i}!)^{m_{i}}}.$
The coefficient from $T_{k}$ then matches. ∎
As a consequence, given a Maurer-Cartan element $\mathfrak{mc}$, then $\phi(\mathfrak{mc})$ is again a Maurer-Cartan element. Since $\phi(c+x)=c+\phi(x)$ for $c\in R$, we have $\overline{\phi(\overline{\mathfrak{mc}})}=\overline{\phi(\mathfrak{mc})}$.
Given a Maurer-Cartan element $\mathfrak{mc}$, we have a deformed $BL_{\infty}$ structure $p^{k,l}_{\mathfrak{mc}}$ defined by trees from stacking the forests representing $e^{\mathfrak{mc}}-1$ over $\widehat{p}$, namely:
Figure 5. A component of $p^{1,7}_{\mathfrak{mc}}$
In a more formal way, we have
(2.6) $p_{\mathfrak{mc}}^{k,l}(v_{1}\ldots v_{k})=\pi_{1,l}\circ\widehat{p}(v_{1}\odot\ldots\odot v_{k}\odot e^{\mathfrak{mc}})$
where $\pi_{1,l}$ is the completed version of the projection from $EV$ to $S^{l}V\subset SV\subset EV$. Since $\mathfrak{mc}$ has degree 00, it does not matter where to put the Maurer-Cartan element in Figure 5. The same construction holds for any $a\in\overline{SV\otimes_{{\bm{k}}}R}$ of degree 00, and we have the following.
Lemma 2.18.
For $s\in\overline{EV\otimes_{{\bm{k}}}R}$ and $a\in\overline{SV\otimes_{{\bm{k}}}R}$ such that $\deg(a)=0$, we have
$\widehat{p}(s\odot e^{a})=\widehat{p}_{a}(s)\odot e^{a}+(-1)^{|s|}s\odot\widehat{p}(e^{a}-1),$
assuming $s$ is of pure degree.
Proof.
The proof is again by examining that both sides correspond to the same linear combination of forests. The left-hand side can be interpreted as summing over the following forests (with coefficient from $e^{a}$) with the top level representing $s$.
While the same forest can be interpreted the other way. Namely, when the $p^{k,l}$ component is glued to the forest representing $s$, then it corresponds to a component of $\widehat{p}_{a}(s)\odot e^{a}$. And when the $p^{k,l}$ component is not glued to $s$, it corresponds to a component of $(-1)^{|s|}s\odot\widehat{p}(e^{a}-1)$. The coefficients match by the same argument of Lemma 2.17. The extra sign of the last term comes from the fact that $|\widehat{p}|=1$. ∎
Then $p^{k,l}_{\mathfrak{mc}}$ is the structural map under the change of coordinates by $\exp_{\mathfrak{mc}}$. More precisely, we have the following.
Proposition 2.19.
For a Maurer-Cartan element $\mathfrak{mc}$ and a filtered completed $BL_{\infty}$ algebra $p^{k,l}$, we have $p^{k,l}_{\mathfrak{mc}}$ forms a filtered completed $BL_{\infty}$ algebra. Moreover, the deformed structure does not depend on the constant term of $\mathfrak{mc}$.
Proof.
By Lemma 2.18, $p^{k,l}_{\mathfrak{mc}}(v_{1},\ldots,v_{k})=\pi_{1,l}\circ\exp_{-\mathfrak{mc}}\circ\widehat{p}\circ\exp_{\mathfrak{mc}}(v_{1}\odot\ldots\odot v_{k})$ and $\widehat{p}_{\mathfrak{mc}}=\exp_{-\mathfrak{mc}}\circ\widehat{p}\circ\exp_{\mathfrak{mc}}$. Therefore $\widehat{p}_{\mathfrak{mc}}\circ\widehat{p}_{\mathfrak{mc}}(s)=\exp_{-\mathfrak{mc}}\circ\widehat{p}\circ\widehat{p}\circ\exp_{\mathfrak{mc}}(s)=0$. That $p^{k,l}_{\mathfrak{mc}}$ does not depend on the constant term of $\mathfrak{mc}$ follows easily from definitions. ∎
Another instant corollary of Lemma 2.18 is the following.
Proposition 2.20.
Let $\mathfrak{mc}$ be a Maurer-Cartan element, then $\exp_{\mathfrak{mc}}$ defines a chain map from $(\overline{EV\otimes_{{\bm{k}}}R},\widehat{p}_{\mathfrak{mc}})$ to $(\overline{EV\otimes_{{\bm{k}}}R},\widehat{p})$.
However, since $\exp_{\mathfrak{mc}}$ does not preserve the sentence length filtration $\overline{E^{k}V\otimes_{{\bm{k}}}R}\subset\overline{EV\otimes_{{\bm{k}}}R}$, we do not have that the torsion of $p$ is smaller than that of $p_{\mathfrak{mc}}$.
Definition 2.21.
Let $p_{\bullet}$ be a filtered completed pointed map on a completed $BL_{\infty}$ algebra $(V,p)$ and $\mathfrak{mc}$ a Maurer-Cartan element. We say $\mathfrak{mc}_{\bullet}\in\overline{SV\otimes_{{\bm{k}}}R}$ of degree $|p_{\bullet}|+1$ with a positive filtration degree is a pointed Maurer-Cartan element if $\widehat{p}_{\bullet}(e^{\mathfrak{mc}}-1)=\widehat{p}_{\mathfrak{mc}}(\mathfrak{mc}_{\bullet})$.
We want to use $\exp_{\mathfrak{mc}}$ as a change of coordinates to deform $\widehat{p}_{\bullet}$. The analog of (2.6) holds for $p_{\bullet}$ as well. However, $\widehat{p}_{\bullet}(e^{\mathfrak{mc}}-1)$ may not vanish, we will explain below that a pointed Maurer-Cartan element is to correct this error in the deformed pointed map. In the context of SFT, when the Maurer-Cartan element arises from a strong cobordism $W$ and the pointed map is defined by counting holomorphic subject to a constraint in a closed chain $C$ in the negative boundary $\partial_{-}W$, the pointed Maurer-Cartan element is defined by counting holomorphic curves without positive puncture in the cobordism subject to a closed chain $D$ in $(W,\partial W)$ such that the boundary of $D$ in $\partial_{-}W$ is $C$, see §3.3 for details.
Proposition 2.22.
Let $\phi$ be a filtered completed $BL_{\infty}$ morphism from $(V,p)$ to $(V^{\prime},q)$ and $\mathfrak{mc}$ a Maurer-Cartan element of $V$.
  1. (1)
    We have a deformed filtered completed $BL_{\infty}$ morphism $\phi_{\mathfrak{mc}}$ from $(V,p_{\mathfrak{mc}}=p_{\overline{\mathfrak{mc}}})$ to $(V^{\prime},q_{\phi\circ\mathfrak{mc}}=q_{\overline{\phi\circ\mathfrak{mc}}})$.
  2. (2)
    Given a filtered completed pointed map $p_{\bullet}$ for $V$ and a pointed Maurer-Cartan element $\mathfrak{mc}_{\bullet}$, we have a deformed filtered pointed map $p_{\bullet,\mathfrak{mc},\mathfrak{mc}_{\bullet}}$ for $p_{\mathfrak{mc}}$.
  3. (3)
    If, in addition, we have a pointed map $q_{\bullet}$ for $V^{\prime}$ compatible with $\phi$, then $p_{\bullet,\mathfrak{mc},\mathfrak{mc}_{\bullet}},q_{\bullet,\phi(\mathfrak{mc}),\mathfrak{mc}^{\prime}_{\bullet}},\phi_{\mathfrak{mc}}$ are compatible, where $\mathfrak{mc}^{\prime}_{\bullet}$ (depends on $\phi,\phi_{\bullet},\mathfrak{mc},\mathfrak{mc}_{\bullet}$ and is defined in the proof) is a pointed Maurer-Cartan element.
Proof.
$\phi_{\mathfrak{mc}}$ is defined similarly to $p_{\mathfrak{mc}}$, i.e. by the following glued tree:
Figure 6. A component of $\phi^{1,6}_{\mathfrak{mc}}$
They do not depend on the constant terms of $\mathfrak{mc}$. The formal definition is similar to (2.6), i.e.
$\phi^{k,l}(v_{1}\ldots v_{k})=\pi_{1,l}\circ\widehat{\phi}\circ\exp_{\mathfrak{mc}}(v_{1}\odot\ldots\odot v_{k}).$
Similar to Lemma 2.18, we have
$\widehat{\phi}\circ\exp_{\mathfrak{mc}}=\exp_{\phi\circ\mathfrak{mc}}\circ\widehat{\phi}_{\mathfrak{mc}}.$
Here, since the composition with $\widehat{\phi}$ glues every vertex of the input forest, we do not have the extra term in (2.6). Therefore we have
$\widehat{\phi}_{\mathfrak{mc}}=\exp_{-\phi\circ\mathfrak{mc}}\circ\widehat{\phi}\circ\exp_{\mathfrak{mc}}.$
Hence from $\widehat{\phi}\circ\widehat{p}=\widehat{q}\circ\widehat{\phi}$ we have
$\widehat{\phi}_{\mathfrak{mc}}\circ\widehat{p}_{\mathfrak{mc}}=\widehat{q}_{\phi\circ\mathfrak{mc}}\circ\phi_{\mathfrak{mc}},$
that is (1) holds.
The deformed version of $p_{\bullet}$ is defined by trees from gluing $e^{\mathfrak{mc}}-1$ to $\widehat{p}_{\bullet}$ and gluing $\mathfrak{mc}_{\bullet}\odot e^{\mathfrak{mc}}$ to $\widehat{p}$. In terms of formula,
$p^{k,l}_{\bullet,\mathfrak{mc},\mathfrak{mc}_{\bullet}}(v_{1}\ldots v_{k}):=\pi_{1,l}\left(p_{\bullet}\circ\exp_{\mathfrak{mc}}(v_{1}\odot\ldots\odot v_{k})-\widehat{p}\circ\exp_{\mathfrak{mc}}(\mathfrak{mc}_{\bullet}\odot v_{1}\odot\ldots\odot v_{k})\right).$
Similar to the proof of Lemma 2.18, we can get
$\exp_{\mathfrak{mc}}\circ\widehat{p}_{\bullet,\mathfrak{mc},\mathfrak{mc}_{\bullet}}(s)=\widehat{p}_{\bullet}\circ\exp_{\mathfrak{mc}}(s)-\widehat{p}\circ\exp_{\mathfrak{mc}}(\mathfrak{mc}_{\bullet}\odot s)+(-1)^{|\mathfrak{mc}_{\bullet}|}\mathfrak{mc}_{\bullet}\odot\widehat{p}\circ\exp_{\mathfrak{mc}}(s)$
Here $\widehat{p}\circ\exp_{\mathfrak{mc}}(\mathfrak{mc}_{\bullet}\odot s)$ counts forests of gluing $\widehat{p}$ to $\exp_{\mathfrak{mc}}(\mathfrak{mc}_{\bullet}\odot s)$. However, in $p^{k,l}_{\bullet,\mathfrak{mc},\mathfrak{mc}_{\bullet}}$, we only count those with the $p$-component glued the $\mathfrak{mc}_{\bullet}$ component by definition. The extra term $(-1)^{|\mathfrak{mc}_{\bullet}|}\mathfrak{mc}_{\bullet}\odot\widehat{p}\circ\exp_{\mathfrak{mc}}(s)$ is to compensate the over-count. Therefore we have
$\widehat{p}_{\mathfrak{mc}}\circ\widehat{p}_{\bullet,\mathfrak{mc},\mathfrak{mc}_{\bullet}}(s)$ $=$ $\exp_{-\mathfrak{mc}}\circ\widehat{p}\circ\widehat{p}_{\bullet}\circ\exp_{\mathfrak{mc}}(s)+(-1)^{|\mathfrak{mc}_{\bullet}|}\exp_{-\mathfrak{mc}}\circ\widehat{p}(\mathfrak{mc}_{\bullet}\odot\widehat{p}\circ\exp_{\mathfrak{mc}}(s))$
$=$ $(-1)^{|p_{\bullet}|}\exp_{-\mathfrak{mc}}\circ\widehat{p}_{\bullet}\circ\widehat{p}\circ\exp_{\mathfrak{mc}}(s)+(-1)^{|\mathfrak{mc}_{\bullet}|}\exp_{-\mathfrak{mc}}\circ\widehat{p}\circ\exp_{\mathfrak{mc}}(\mathfrak{mc}_{\bullet}\odot\widehat{p}_{\mathfrak{mc}}(s))$
and
$\widehat{p}_{\bullet,\mathfrak{mc},\mathfrak{mc}_{\bullet}}\circ\widehat{p}_{\mathfrak{mc}}(s)$ $=$ $\exp_{-\mathfrak{mc}}\circ\widehat{p_{\bullet}}\circ\widehat{p}\circ\exp_{\mathfrak{mc}}(s)-\exp_{-\mathfrak{mc}}\circ\widehat{p}\circ\exp_{\mathfrak{mc}}(\mathfrak{mc}_{\bullet}\odot\widehat{p}_{\mathfrak{mc}}(s))$
As $|\mathfrak{mc}_{\bullet}|=|p_{\bullet}|+1$, we have $p_{\bullet,\mathfrak{mc},\mathfrak{mc}_{\bullet}}$ is a pointed map of degree $|p_{\bullet}|$ for $p_{\mathfrak{mc}}$. This proves (2).
The deformed version of $\phi_{\bullet}$ is defined by trees from gluing $e^{\mathfrak{mc}}-1$ to $\widehat{\phi}_{\bullet}$ and gluing $\mathfrak{mc}_{\bullet}\odot e^{\mathfrak{mc}}$ to $\widehat{\phi}$, the pointed Maurer-Cartan element $\mathfrak{mc}^{\prime}_{\bullet}$ is defined by trees (without input vertices) from gluing $e^{\mathfrak{mc}}-1$ and $\widehat{\phi}_{\bullet}$ and gluing $\mathfrak{mc}_{\bullet}\odot e^{\mathfrak{mc}}$ to $\widehat{\phi}$. In terms of formula, we have
$\mathfrak{mc}^{\prime}_{\bullet}=\pi_{1}(\widehat{\phi}_{\bullet}(e^{\mathfrak{mc}})+\widehat{\phi}(\mathfrak{mc}_{\bullet}\odot e^{\mathfrak{mc}}))$
where $\pi_{1}$ is the projection from $\overline{EV\otimes_{{\bm{k}}}R}$ to $\overline{SV\otimes_{{\bm{k}}}R}$ and the deformed version of $\phi_{\bullet}$ is defined by
$\phi_{\bullet,\mathfrak{mc},\mathfrak{mc}_{\bullet}}^{k,l}(v_{1}\ldots v_{k})=\pi_{1,l}\left(\widehat{\phi}_{\bullet}\circ\exp_{\mathfrak{mc}}(v_{1}\odot\ldots\odot v_{k})+\widehat{\phi}\circ\exp_{\mathfrak{mc}}(\mathfrak{mc}_{\bullet}\odot v_{1}\odot\ldots\odot v_{k})\right).$
The desired algebraic properties follow from a similar argument as before. ∎
3. Rational symplectic field theory
Unless specified otherwise, $(Y,\alpha)$ is a strict compact contact manifold such that the contact form $\alpha$ is non-degenerate throughout this section. By exact strict cobordism from $(Y_{-},\alpha_{-})$ to $(Y_{+},\alpha_{+})$, we mean a Liouville domain $(X,\lambda)$ such that $(Y_{-},\alpha_{-})$ is the concave boundary $(\partial_{-}X,\lambda)$ and $(Y_{+},\alpha_{+})$ is the convex boundary $(\partial_{+}X,\lambda)$ as strict contact manifolds.
3.1. RSFT of contact manifolds
3.1.1. RSFT as a $BL_{\infty}$ algebra
To assign a $BL_{\infty}$ algebra to a strict contact manifold $(Y,\alpha)$, let $V_{\alpha}$ denote the free ${\mathbb{Q}}$-module generated by formal variables $q_{\gamma}$ for each (good) orbit $\gamma$ of $(Y,\alpha)$ and $\Lambda$ the Novikov field over ${\mathbb{Q}}$. We take the grading $|q_{\gamma}|=\mu_{CZ}(\gamma)+n-3$. We define $p^{k,l}:S^{k}V_{\alpha}\otimes_{{\mathbb{Q}}}\Lambda\rightarrow S^{l}V_{\alpha}\otimes_{{\mathbb{Q}}}\Lambda$ by
(3.1) $p^{k,l}(q^{\Gamma^{+}})=\sum_{A,|[\Gamma^{-}]|=l}\#\overline{{\mathcal{M}}}_{Y,A}(\Gamma^{+},\Gamma^{-})\frac{T^{\int_{A}{\rm d}\alpha}}{\mu_{\Gamma^{-}}\kappa_{\Gamma^{-}}}q^{\Gamma^{-}}.$
[…]
  1. (1)
    By setting $T=1$, $p^{k,l}$ defines a $BL_{\infty}$ algebra on the ${\mathbb{Q}}$-vector space $V_{\alpha}$, which was considered in [31, §3]. We use $\operatorname{RSFT}(Y)$ to denote this $BL_{\infty}$ algebra for simplicity.
  2. (2)
    $p^{k,l}$ in (3.1) defines a $BL_{\infty}$ algebra on the $\Lambda$-vectord space $V_{\alpha}\otimes_{{\mathbb{Q}}}\Lambda$, which we referred to as the elementary version over $\Lambda$. We use $\operatorname{RSFT}_{\Lambda}(Y)$ to denote this $BL_{\infty}$ algebra.
  3. (3)
    Finally, $p^{k,l}$ defines a filtered completed $BL_{\infty}$ algebra as in Definition 2.15. We use $\overline{\operatorname{RSFT}}(Y)$ to denote this $BL_{\infty}$ algebra.
To be more precise, those $BL_{\infty}$ algebras depend on the choice of contact form $\alpha$ as well as auxiliary choices in the virtual machinery. However, the homotopy type of them is expected to be independent of those extra choices.
Note that if $u\in\overline{{\mathcal{M}}}_{Y,A}(\Gamma_{+},\Gamma_{-})$, then we have
[…]
As a consequence, (3.1) is a finite sum, hence $\overline{\operatorname{RSFT}}(Y)$ is defined. At the $BL_{\infty}$ algebra level, those three versions carry the same amount of information, as $\operatorname{RSFT}_{\Lambda}$ is a restriction of $\overline{\operatorname{RSFT}}$, and by setting $T=1$ in $\operatorname{RSFT}_{\Lambda}$, we can get $\operatorname{RSFT}$. On the other hand, since the ${\rm d}\alpha$ energy of a curve is determined by its asymptotics, we can recover $\operatorname{RSFT}_{\Lambda}$ and $\overline{\operatorname{RSFT}}$.
3.1.2. Functoriality
Given a strict exact cobordism $X$ from $Y_{-}$ to $Y_{+}$, we define a $BL_{\infty}$ morphism from $\operatorname{RSFT}(Y_{+})$ to $\operatorname{RSFT}(Y_{-})$ (similarly for the other two versions as well) by (setting $T=1$)
(3.2) $\phi^{k,l}(q^{\Gamma^{+}})=\sum_{A,|[\Gamma^{-}]|=l}\#\overline{{\mathcal{M}}}_{X,A}(\Gamma^{+},\Gamma^{-})\frac{T^{\int_{A}{\rm d}\lambda}}{\mu_{\Gamma^{-}}\kappa_{\Gamma^{-}}}q^{\Gamma^{-}},$
[…]
3.1.3. Pointed morphisms
If we fix a point $o$ in $Y$, consider the compactified moduli space ¯$\overline{{\mathcal{M}}}_{Y,A,o}(\Gamma^{+},\Gamma^{-})$ of rational holomorphic curves with one interior marked point in the symplectization modulo automorphisms, where the marked point is required to be mapped to $(0,o)\in\mathbb{R}\times Y$, in homology class $A$ and with asymptotics $\Gamma^{\pm}$. We then define a pointed map $p_{\bullet_{o}}$ by
(3.3) $p^{k,l}_{\bullet_{o}}(q^{\Gamma^{+}})=\sum_{A,|[\Gamma^{-}]|=l}\#\overline{{\mathcal{M}}}_{Y,A,o}(\Gamma^{+},\Gamma^{-})\frac{T^{\int_{A}{\rm d}\alpha}}{\mu_{\Gamma^{-}}\kappa_{\Gamma^{-}}}q^{\Gamma^{-}}.$
Then the boundary of the $1$-dimensional moduli spaces ¯$\overline{{\mathcal{M}}}_{Y,A,o}(\Gamma^{+},\Gamma^{-})$ yields that $\{p^{k,l}_{\bullet_{o}}\}$ is a pointed map of degree 00 for all three versions of rational SFT.
For a strict exact cobordism $X$ from $Y_{-}$ to $Y_{+}$, we choose a path $\gamma$ from $o_{-}\in Y_{-}$ to $o_{+}\in Y_{+}$. We complete the path $\gamma$ to a proper path $\widehat{\gamma}$ in the Liouville completion $\widehat{X}$ by constants (in $Y_{\pm}$) in the cylindrical ends. Then the pointed morphisms $p_{\bullet_{o_{+}}},p_{\bullet_{o_{-}}}$ determined by $o_{-},o_{+}$ and the $BL_{\infty}$ morphism $\phi$ are compatible, with the map $\phi_{\bullet}$ given by
(3.4) $\phi^{k,l}_{\bullet}(q^{\Gamma^{+}})=\sum_{A,|[\Gamma^{-}]|=l}\#\overline{{\mathcal{M}}}_{X,A,\gamma}(\Gamma^{+},\Gamma^{-})\frac{T^{\int_{A}{\rm d}\lambda}}{\mu_{\Gamma^{-}}\kappa_{\Gamma^{-}}}q^{\Gamma^{-}}.$
Here ¯$\overline{{\mathcal{M}}}_{X,A,\gamma}(\Gamma^{+},\Gamma^{-})$ is the moduli space of rational curves with one interior marked point in the completed cobordism $\widehat{X}$ modulo automorphism, in homology class $A$ and with asymptotics $\Gamma^{\pm}$, where the marked point is required to go through $\widehat{\gamma}$.
3.2. Realizing RSFT using Pardons’s VFC
Theorem 3.1 ([31, Theorem 3.10]).
Let $(Y,\alpha)$ be a strict contact manifold with a non-degenerate contact form, then we have the following.
  1. (1)
    There exists a non-empty set of auxiliary data $\Theta$, such that for each $\theta\in\Theta$ we have a $BL_{\infty}$ algebra $p_{\theta}$ on $V_{\alpha}\otimes_{{\mathbb{Q}}}\Lambda$, which induces a $BL_{\infty}$ algebra on $V_{\alpha}$ by setting $T=1$ and a filtered completed $BL_{\infty}$ algebra.
  2. (2)
    For any point $o\in Y$, there exists a set of auxiliary data $\Theta_{o}$ with a surjective map $\Theta_{o}\to\Theta$, such that for any $\theta_{o}\in\Theta_{o}$, we have a pointed map $p_{\bullet,\theta_{o}}$ for $p_{\theta}$, where $\theta$ is the image of $\theta_{o}$ in $\Theta_{o}\to\Theta$.
  3. (3)
    Assume there is a strict exact cobordism $X$ from $(Y^{\prime},\alpha^{\prime})$ to $(Y,\alpha)$. Let $\Theta,\Theta^{\prime}$ be the sets of auxiliary data for $\alpha,\alpha^{\prime}$, then there exists a set of auxiliary data $\Xi$ with a surjective map $\Xi\to\Theta\times\Theta^{\prime}$, such that for $\xi\in\Xi$, there is a $BL_{\infty}$ morphism $\phi_{\xi}$ from $(V_{\alpha}\otimes_{{\mathbb{Q}}}\Lambda,p_{\theta})$ to $(V_{\alpha^{\prime}}\otimes_{{\mathbb{Q}}}\Lambda,p_{\theta^{\prime}})$, where $(\theta,\theta^{\prime})$ is the image of $\xi$ under $\Xi\to\Theta\times\Theta^{\prime}$.
  4. (4)
    Assume in addition that we fix a point $o^{\prime}\in Y^{\prime}$ that is in the same component of $o$ in $X$. Then for any compatible auxiliary data $\theta,\theta^{\prime},\theta_{o},\theta_{o^{\prime}},\xi$, we have $p_{\bullet,\theta_{o}},p_{\bullet,\theta_{o^{\prime}}},\phi_{\xi}$ are compatible.
  5. (5)
    For compatible auxiliary data $\theta,\theta_{o}$, there exists compatible auxiliary data $k\theta,k\theta_{o}$ for $(Y,k\alpha)$ for $k\in\mathbb{R}_{+}$, such that $p_{k\theta},p_{\bullet,k\theta_{o}}$ are identified with $p_{\theta},p_{\bullet,\theta_{o}}$ by the canonical identification between $V_{\alpha}$ and $V_{k\alpha}$ and the isomorphism $\Lambda\to\Lambda$ defined by $T\mapsto T^{1/k}$.
  6. (6)
    A strict strong semi-filling, i.e. a strong filling with $Y$ as a boundary component, gives rise to a $BL_{\infty}$ augmentation for the elementary version over $\Lambda$.
Remark 3.2.
We can assign weights to generators $q_{\gamma}$ of $V_{\alpha}$ by $\int\gamma^{*}\alpha$ and weight $1$ to the formal variable $T$ is the Novikov field. Then all of $p,p_{\bullet},\phi,\phi_{\bullet}$ preserve the weight in $\operatorname{RSFT}_{\Lambda}(Y)$ and $\overline{\operatorname{RSFT}}(Y)$. Note that any element in $EV_{\alpha}\otimes_{{\mathbb{Q}}}\Lambda$ or $\overline{EV_{\alpha}\otimes_{{\mathbb{Q}}}\Lambda}$ (the completion) with pure weight is contained in $\overline{S}^{k}S^{m}V_{\alpha}$ for some $k,m$ with coefficients in $\{\sum_{i=1}^{N}a_{i}T^{t_{i}}|N\in{\mathbb{N}}_{+},a_{i}\in{\mathbb{Q}}\}\subset\Lambda$, as the period $\int\gamma^{*}\alpha$ of a Reeb orbit $\gamma$ is strictly positive and below any period there are only finite Reeb orbits.
3.3. Maurer-Cartan Elements
Let $(X,\omega)$ be a strict strong cobordism from $(Y_{-},\alpha_{-})$ to $(Y_{+},\alpha_{+})$, it induces a Maurer-Cartan element:
$\mathfrak{mc}=\sum_{A,[\Gamma_{-}]}\#\overline{{\mathcal{M}}}_{X,A}(\emptyset,\Gamma_{-})\frac{T^{\int_{A}\overline{\omega}}}{\mu_{\Gamma_{-}}\kappa_{\Gamma_{-}}}q^{\Gamma^{-}}$
where $\overline{\omega}$ defined on $\widehat{X}$ is $\omega$ on $X$ and ${\rm d}\alpha_{-},{\rm d}\alpha_{+}$ on the cylindrical ends of $X$. Even though $\overline{\omega}$ is not smooth on $\widehat{X}$, it is easy to show that $\int_{A}\overline{\omega}$ only depends on the relative homology class by Stokes’ theorem. If $u\in\overline{{\mathcal{M}}}_{X,A}(\emptyset,\Gamma_{-})$, we have $0<\int u^{*}\overline{\omega}=\int_{A}\overline{\omega}$, which is called the $\omega$-energy in [3, §6.1]. Then the Hofer energy of $u$ is bounded above by $2\int u^{*}\overline{\omega}$, hence the SFT compactness [3] implies that $\mathfrak{mc}$ is in the completion $\overline{SV_{\alpha_{-}}\otimes_{{\mathbb{Q}}}\Lambda}$ of a positive filtration degree assuming the counting is taken care of by some virtual technique. We can also define a map $\phi^{k,l}:\overline{S^{k}V_{\alpha_{+}}\otimes_{{\mathbb{Q}}}\Lambda}$ to $\overline{S^{l}V_{\alpha_{-}}\otimes_{{\mathbb{Q}}}\Lambda}$ for $k\geq 1$ by
$\phi^{k,l}(q^{\Gamma^{+}})=\sum_{A,|[\Gamma^{-}]|=l}\#\overline{{\mathcal{M}}}_{X,A}(\Gamma^{+},\Gamma^{-})\frac{T^{\int_{A}\overline{\omega}}}{\mu_{\Gamma^{-}}\kappa_{\Gamma^{-}}}q^{\Gamma^{-}}.$
It is well-defined in the completion as the Hofer energy of $u\in\overline{{\mathcal{M}}}_{X,A}(\Gamma^{+},\Gamma^{-})$ is bounded by $2\int u^{*}\overline{\omega}+\sum_{\gamma\in\Gamma_{+}}\int\gamma^{*}\alpha_{+}$.
Similarly, for a path $\gamma$ in $X$ with ends $o_{-}$ in $Y_{-}$ and $o_{+}$ in $Y_{+}$, we have
$\mathfrak{mc}_{\bullet}=\sum_{A,[\Gamma_{-}]}\#\overline{{\mathcal{M}}}_{X,A,\gamma}(\emptyset,\Gamma_{-})\frac{T^{\int_{A}\overline{\omega}}}{\mu_{\Gamma_{-}}\kappa_{\Gamma_{-}}}q^{\Gamma^{-}}$
is a pointed Maurer-Cartan element for the pointed map induced from $o_{-}$. It is helpful to point out that a curve in ¯$\overline{{\mathcal{M}}}_{X,A,\gamma}(\emptyset,\Gamma_{-})$ can not pass through the ray in $Y_{+}\times\mathbb{R}_{+}$ from $o_{+}$ by the maximum principle. Then the boundary configuration of ¯$\overline{{\mathcal{M}}}_{X,A,\gamma}(\emptyset,\Gamma_{-})$ implies the algebraic property for a pointed Maurer-Cartan element. Finally $\phi_{\bullet}$ from $X$ is defined similarly.
As the Maurer-Cartan deformed theory has a description by trees, the following proposition follows from the same argument of Theorem 3.1.
Proposition 3.3.
Assume $X$ a strict strong cobordism from $(Y_{-},\alpha_{-})$ to $(Y_{+},\alpha_{+})$, then there exists a set of auxiliary data $\Xi$ and a surjective map $\Xi\to\Theta_{-}\times\Theta_{+}$, such that for each $\xi\in\Xi$, which is mapped to $(\theta_{-},\theta_{+})\in\Theta_{-}\times\Theta_{+}$, we have a Maurer-Cartan element $\mathfrak{mc}_{\xi}$ of $p_{\theta_{-}}$ and a filtered completed $BL_{\infty}$ morphism $\phi_{\xi}$ from $p_{\theta_{+}}$ to $p_{\theta_{-},\mathfrak{mc}_{\xi}}$. Assume $o_{+},o_{-}$ be two points in $Y_{+},Y_{-}$ that are in the same component of $X$, then $\phi_{\xi},p_{\bullet,\theta_{o_{+}}},p_{\bullet,\theta_{o_{-}},\mathfrak{mc}_{\xi},\mathfrak{mc}_{\bullet}}$ are compatible with a pointed Maurer-Cartan element $\mathfrak{mc}_{\bullet}$.
4. Invariants
4.1. Algebraic planar torsion
As explained in [31], while setting up the theory requires the choice of contact forms and auxiliary data coming from virtual techniques, the following is independent of choices:
Definition 4.1 ([31, Proposition 3.11]).
For a contact manifold $(Y,\xi)$, we define its algebraic planar torsion $\operatorname{APT}(Y)$ to be the torsion of the $BL_{\infty}$ algebra $\operatorname{RSFT}(Y,\alpha,\theta)$, for any choice of contact form $\alpha$ and auxiliary choice $\theta$ in Theorem 3.1.
Using the weight in Remark 3.2, algebraic planar torsion is independent of the choice of coefficients, i.e. ${\mathbb{Q}},\Lambda$ and the completed version over $\Lambda$. The finiteness of $\operatorname{APT}$ is an obstruction to $BL_{\infty}$ augmentations hence implies no strong filling (by (6) of Theorem 3.1) and $\operatorname{APT}$ is non-decreasing in an exact cobordism. We refer to [31, §3.5] for properties that are enjoyed by this invariant as well as their proofs.
Proof of Theorem 1.1.
By assumption, $\operatorname{APT}(Y_{+})<+\infty$, so there exists an element $v\in E^{k}V_{\alpha_{+}}\otimes_{{\mathbb{Q}}}{\Lambda}\subset\overline{EV_{\alpha_{+}}\otimes_{{\mathbb{Q}}}\Lambda}$ of weight 00 (Remark 3.2), such that $\widehat{p}_{+}(v)=1$. In particular, the Novikov coefficients of $v$ all have negative powers. By Propositions 2.20 and 3.3, we have filtered chain maps
$(\overline{EV_{\alpha_{+}}\otimes_{{\mathbb{Q}}}\Lambda},\widehat{p}_{+})\stackrel{{\scriptstyle\widehat{\phi}}}{{\longrightarrow}}(\overline{EV_{\alpha_{-}}\otimes_{{\mathbb{Q}}}\Lambda},\widehat{p}_{-,\mathfrak{mc}})\stackrel{{\scriptstyle\exp_{\mathfrak{mc}}}}{{\longrightarrow}}(\overline{EV_{\alpha_{-}}\otimes_{{\mathbb{Q}}}\Lambda},\widehat{p}_{-}).$
As a consequence, we have $\widehat{p}_{-}\circ\exp_{\mathfrak{mc}}\circ\widehat{\phi}(v)=\exp_{\mathfrak{mc}}(1)=1\odot e^{\mathfrak{mc}}$ for $1\in SV\subset\overline{EV\otimes_{{\bm{k}}}R}$. Let $v_{0}$ be the weight zero part of $\exp_{\mathfrak{mc}}\circ\widehat{\phi}(v)$, then $\widehat{p}_{-}(v_{0})=1$ as the weight zero part of $1\odot e^{\mathfrak{mc}}$ is $1$. Then there must be $m\in{\mathbb{N}}$, such that $v_{0}\in E^{m}V_{\alpha_{-}}\otimes_{{\mathbb{Q}}}\Lambda$ by Remark 3.2, i.e. $\operatorname{APT}(Y_{-})<\infty$. ∎
Example 4.2.
It is expected that containing a bordered Legendrian open book (bLob) [32, Section 3] would imply that the contact homology vanishes, i.e. $\operatorname{APT}=0$. Assuming this is established, the cobordism in [32], along with Theorem 1.1, yields another proof that the algebraic planar torsion of contact manifolds with higher dimensional Giroux torsion [32] is finite. Such results were established by the first author in [30] using a more direct approach with explicit upper bounds. The functorial approach using Theorem 1.1 can be refined in a similar way to Corollaries 1.4 and 1.6 to obtain upper bounds if we know more precisely about how finite algebraic torsion is formed for the convex boundary.
Remark 4.3.
In [25, Question 2], Latschev and Wendl asked for the version of Theorem 1.1 for algebraic torsions [25, Definition 1.1]. This follows from the same proof using the weight zero part of the deformed functoriality in the Maurer-Cartan formalism for $BV_{\infty}$ algebras, assuming Theorem 3.1 and Proposition 3.3 hold for the $BV_{\infty}$ formalism of the full SFT in [11, 25].
4.2. Planarity
If $\operatorname{RSFT}_{{\mathbb{Q}}}(Y)=(V_{\alpha},p)$ has a $BL_{\infty}$ augmentation $\epsilon$ over ${\mathbb{Q}}$, then $\operatorname{APT}(Y)$ is $\infty$, but we may still define a meaningful invariant as follows. Given a point $o\in Y$, we consider the associated pointed morphism $p_{\bullet}$, and the corresponding order $O(V_{\alpha},\epsilon,p_{\bullet})$. Let $\operatorname{Aug}_{{\mathbb{Q}}}(V_{\alpha})$ denote the set of $BL_{\infty}$ augmentations of $V_{\alpha}$ over ${\mathbb{Q}}$. Again, the following definitions is independent on all auxiliary choices [31]:
Definition 4.4 ([31, Definition 3.21]).
For a contact manifold $(Y,\xi)$, we define its planarity as
$\operatorname{P}(Y):=\max\left\{O(V_{\alpha},\epsilon,p_{\bullet})\left|\forall\epsilon\in\operatorname{Aug}_{{\mathbb{Q}}}(V_{\alpha}),o\in Y,\theta_{o}\in\Theta_{o}\right.\right\},$
where the maximum of an empty set is defined to be zero. Here, $\alpha$ is any choice of contact form.
The basic property of $\operatorname{P}$ is that $\operatorname{P}(Y_{+})\leq\operatorname{P}(Y_{+})$ if there is an exact cobordism from $Y_{-}$ to $Y_{+}$. We again refer to [31, §3.5] for the properties concerning this invariant. Similarly, we can define $\operatorname{P}_{\Lambda}$ and $\overline{\operatorname{P}}_{\Lambda}$ using the elementary version over $\Lambda$ and the filtered completed version of the rational SFT, where in both cases we require the augmentation is filered, i.e. preserves the filtration. Since an augmentation over ${\mathbb{Q}}$ induces a filtered augmentation over $\Lambda$ (which preserves the weight in Remark 3.2), and a filtered augmentation over $\Lambda$ induces a filtered completed augmentation over $\Lambda$, we have that $\operatorname{P}\leq\operatorname{P}_{\Lambda}\leq\overline{\operatorname{P}}_{\Lambda}$. All of them enjoy the same functorial property w.r.t. exact cobordisms by the argument in [31, §3.5].
Proof of Theorem 1.2.
Assume there is a connected strong filling $W$ of $Y\sqcup Y^{\prime}$, which induces a $\Lambda$-augmentation $\epsilon$ of $\operatorname{RSFT}_{\Lambda}(Y)$. By taking a point $o\in Y$, we get a pointed map $p_{\bullet}$. If we consider curves in $\widehat{W}$ with positive punctures only asymptotic to Reeb orbits on $Y$ and with one marked point going through the completion of a ray $\gamma$ starting from $o$ and shooting out to infinity on the cylindrical end of $Y^{\prime}$, we get a structural map $\epsilon_{\bullet}$, which is a special case of $\phi_{\bullet}$, i.e. $(\operatorname{RSFT}(Y),p,p_{\bullet}),(0,0,0),\epsilon$ are compatible through $\epsilon_{\bullet}$. Even though $\gamma$ is non-compact, the moduli space is compactified as usual as there is no curve with only positive punctures asymptotic to orbits on $Y$ going through the infinite end of $\gamma$ in +×$\mathbb{R}_{+}\times Y^{\prime}$ by the maximum principle. Then by [31, Proposition 2.20], we have $O(\operatorname{RSFT}(Y),\epsilon,p_{\bullet})\geq O(\{0\},0,0)=+\infty$, contradiction! ∎
Proposition 4.5 (Theorem 1.3).
Let $X$ be a strong cobordism from $Y_{-}$ to $Y_{+}$. Assume there are auxiliary data $\alpha_{+},\theta_{+}$ for $Y_{+}$, such that the following holds:
  1. (1)
    There exist Reeb orbits $\gamma_{1},\ldots,\gamma_{k}$ of $\alpha_{+}$, such that for any subset $S\subset\{1,\ldots,k\}$, we have $p^{|S|,l}(\prod_{i\in S}q_{\gamma_{i}})=0$ for any $l\geq 0$.
  2. (2)
    For any point $o$ with associated pointed map $p_{\bullet}$, we have $p^{k,0}_{\bullet}(\prod_{i=1}^{k}q_{\gamma_{i}})\neq 0$ and $p^{|S|,l}_{\bullet}(\prod_{i\in S}q_{\gamma_{i}})=0$ for any $l>0$ and subset $S\subset\{1,\ldots,k\}$.
Those two conditions imply that $\overline{\operatorname{P}}_{\Lambda}(Y_{+})\leq k$ (as the choice of augmentation does not affect). Then $\overline{\operatorname{P}}_{\Lambda}(Y_{-})<+\infty$.
Proof.
Using the weight in Remark 3.2, the operation determined by ¯$\overline{{\mathcal{M}}}_{X,A}(\Gamma_{+},\Gamma_{-})$ will increase the weight by $\int_{A}\overline{\omega}+\sum_{\gamma\in\Gamma_{-}}\int\gamma^{*}\alpha_{-}-\sum_{\gamma\in\Gamma_{+}}\int\gamma^{*}\alpha_{+}$, which is zero whenever $X$ is exact. The functorial property in [31, Proposition 2.20] and Proposition 2.22 for the order uses one level of curves from ¯$\overline{{\mathcal{M}}}_{X,A}(\Gamma_{+},\Gamma_{-})$ and ¯$\overline{{\mathcal{M}}}_{X,A,\gamma}(\Gamma_{+},\Gamma_{-})$ for a path $\gamma\subset X$ connecting two points $o_{\pm}\in Y_{\pm}$ respectively, stacking over or under curves in the symplectization defining $p$ and $p_{\bullet}$, and finally cap off from the bottom using an augmentation $\epsilon$ for the RSFT of $Y_{-}$. Since both $p,p_{\bullet}$ have weight 00, we can only consider those graphs with total weight change 00 (i.e. a curve component in $X$ can modify weight, but the whole level must keep the weight). More precisely, we consider the following (output) element $v$ in $\overline{\overline{S}V_{\alpha_{-}}\otimes_{{\mathbb{Q}}}\Lambda}$:
$q_{\gamma_{1}}$$q_{\gamma_{2}}$$q_{\gamma_{3}}$$q_{\gamma_{4}}$$q_{\gamma_{5}}$$\phi^{1,1}$$\phi^{1,1}$$\phi^{2,3}$$\phi^{1,1}$$\mathfrak{mc}_{2}$$\mathfrak{mc}_{1}$$\epsilon_{2}$$\epsilon_{1}$$\epsilon_{2}$
Figure 7. A component of the output element $v$ in $\overline{S^{4}V_{\alpha_{-}}\otimes_{{\mathbb{Q}}}\Lambda}$ when $k=5$
Namely stacking a level of $\widehat{\phi}$ and $e^{\mathfrak{mc}}$ with total weight change 00 to $q_{\gamma_{1}}\odot\ldots\odot q_{\gamma_{k}}$, capped off with some $\epsilon$, such that the output $v$ is in $\overline{\overline{S}V_{\alpha_{-}}\otimes_{{\mathbb{Q}}}\Lambda}=\overline{\oplus_{i\geq 1}\odot^{i}V_{\alpha_{-}}\otimes_{{\mathbb{Q}}}\Lambda}$. Since the middle level has a total weight change of 00, i.e. the sum of $\int_{A}\overline{\omega}+\sum_{\gamma\in\Gamma_{-}}\int\gamma^{*}\alpha_{-}-\sum_{\gamma\in\Gamma_{+}}\int\gamma^{*}\alpha_{+}$ for curves in the level is zero, and $\int u^{*}\overline{\omega}\geq 0$, the word length of $v$ is at most
$N:=\frac{\displaystyle\sum_{i=1}^{k}\int\gamma_{i}^{*}\alpha_{+}}{\displaystyle\min\left\{\int\gamma^{*}\alpha_{-}\left|\gamma\text{ is a Reeb orbit}\right.\right\}}<\infty.$
We have $\widehat{\ell_{\epsilon}}(v)=0$ by (1) of the assumptions and the quadratic relations from disconnected curves in $X$ with total weight change 00. Similarly, we have ^$\widehat{\ell}_{\bullet,\epsilon}(v)\neq 0$ by (2) of the assumptions and the quadratic relations from disconnected curves in $X$ with a point constraint in $\gamma$ and with total weight change 00. Therefore $\overline{\operatorname{P}}_{\Lambda}(Y_{-})\leq N$. ∎
Remark 4.6.
The conditions in Proposition 4.5 are stronger than having an element $x$ in $S^{k}V_{\alpha_{+}}\otimes_{{\mathbb{Q}}}\Lambda$, such that for any filtered augmentation $\epsilon$ of $\operatorname{RSFT}_{\Lambda}(Y_{+})$ we have ^$\widehat{\ell}_{\epsilon}(x)=0$ and ^$\widehat{\ell}_{\bullet,\epsilon}(x)\neq 0$. Note that in the proof of Proposition 4.5 we use the level of curves in $X$ of total weight change 00. To analyzing breakings in the proof of ^$\widehat{\ell}_{\epsilon}$-closedness of $v$ in Proposition 4.5, breaking at the top-level (i.e. $Y_{+}$) involves only part of ^$\widehat{\ell}_{\phi(\epsilon)}$ for augmentation $\epsilon$ of $Y_{-}$ as components of ^$\widehat{\ell}_{\phi(\epsilon)}$ can change weight ($X$ is not exact and $\epsilon$ can change weight). As a consequence, the algebraic relation ^$\widehat{\ell}_{\epsilon}(x)=0$ above may not yield the closedness of $v$. On the other hand, the conditions in Proposition 4.5 imply that breaking at the top-level vanishes geometrically, similarly for the non-vanishing of ^$\widehat{\ell}_{\epsilon,\bullet}(v)$, where conditions Proposition 4.5 simplify the top level geometrically when we analyze the breakings. The computations of $\operatorname{P}$ on explicit examples in [31] are through establishing conditions in Proposition 4.5 above.
Proof of Corollary 1.4 and Corollary 1.6.
In both cases, the contact manifold has a strong cobordism to a subcritically fillable contact manifold $\partial(V\times{\mathbb{D}})$. In the first case, the cobordism is from adding back $s^{-1}(0)\backslash H$, where the convex boundary is the standard contact $S^{2n-1}=\partial({\mathbb{D}}^{n-1}\times{\mathbb{D}})$. In the latter case, the cobordism is from attaching a round handle in [32, §5], see [7] for details. By [31, Theorem L (3)], we have $\operatorname{P}(\partial(V\times{\mathbb{D}}))=1$ such that conditions in Proposition 4.5 hold for $k=1$ and $\gamma_{1}$ is the simple Reeb orbit responsible for the vanishing of $SH^{*}(V\times{\mathbb{D}})$, i.e. the Reeb orbit over the minimum of the Morse function using the contact form in [39, §2.1]. $\gamma_{1}$ bounds a holomorphic plane $u$ with a point constraint in the symplectization of the convex boundary. In the first case, by [31, Lemma 7.1], it is sufficient to prove the case when $s^{-1}(0)$ is a smooth degree $d$ hypersurface that is transverse to $H$. Note that $s^{-1}(0)\backslash H$ is a symplectic hypersurface, whose intersection with the contact boundary $S^{2n-1}$ is the contact submanifold $\pi^{-1}(s^{-1}(0)\cap H)$, where $\pi:S^{2n-1}\to H$ is the Hopf-fibration. The intersection number of $u$ with the symplectization of $\pi^{-1}(s^{-1}(0)\cap H)$ is $d$. If we view this curve as a boundary component of the moduli space of curves with a point constraint in the cobordism, then its intersection number with $s^{-1}(0)\backslash H$ is $d$. In the following, we will use an almost complex structure making the symplectic hypersurface holomorphic. In view of the proof of Proposition 4.5, we can establish the same proof if we only consider those with middle-level whose intersection number with $s^{-1}(0)\backslash H$ is $d$. Since the complement of $s^{-1}(0)\backslash H$ is exact, the curve contributing to the Maurer-Cartan element must intersect $s^{-1}(0)\backslash H$. As a consequence, the output element $v$ in proposition 4.5 is in $S^{d+1}V_{\alpha_{-}}\otimes_{{\mathbb{Q}}}\Lambda$, i.e. the planarity of the concave boundary is at most $d+1$. The argument for Bourgeois contact manifolds is similar using the symplectic core of the round handle. ∎
Theorem 4.7 ([31, Theorem 7.14]).
Assume $Y$ be the Boothby-Wang contact structure over a symplectic manifold $(X,\omega,J)$, such that there is no $J$-holomorphic spheres in $X$. Then $\operatorname{P}(Y)=\infty$.
Proof of Theorem 1.7.
By [26, Lemma 2.5, 2.12], $Y_{d,g}$ is not cofillable if and only if $d>2g-2$. By Theorem 4.7 and Proposition 4.5, there is no strong cobordism from $Y_{d,g}$ to $(S^{3},\xi_{\mathrm{std}})$ for any $d>0$. ∎
Remark 4.8.
One can also use the classification of strong fillings of $(S^{3},\xi_{\mathrm{std}})$ due to Gromov [21], Eliashberg [14] and McDuff [28] to obstruct a strong cobordism. If there is a strong cobordism from $Y_{d,g}$ to $(S^{3},\xi_{\mathrm{std}})$, then we can glue it with the natural disk bundle filling of $Y_{d,g}$ to get strong filling $W$ of $(S^{3},\xi_{\mathrm{std}})$, which is diffeomorphic
$B^{4}\#\overline{\mathbb{C}\mathbb{P}^{2}}\ldots\#\overline{\mathbb{C}\mathbb{P}^{2}}.$
Assume the zero section $\Sigma_{g}$ of disk bundle filling is represented by $A=\sum k_{i}E_{i}$ in homology. Then we have
$-\sum k^{2}_{i}=A\cdot A=\langle c_{1}(W),A\rangle+2g-2=\sum k_{i}^{2}+2g-2$
which is impossible if $g\geq 1$.
5. Generalized constraints and higher genera invariants
From here onwards, we discuss the addenda to [31].
5.1. Generalized constraints
There are a few directions where one can generalize the constructions in [31]. In the following, we will briefly describe one of such generalizations by considering more general constraints.
5.1.1. General constraints from $Y$
We can pick a closed submanifold of $Y$ or more generally a closed singular chain $C$ of $Y$ to construct a pointed morphism $p_{C}$ by considering rational holomorphic curves in ×$\mathbb{R}\times Y$ with one marked point mapped to $\{0\}\times C$. Then given a $BL_{\infty}$ augmentation $\epsilon$, we have an induced chain map ^$\widehat{\ell}_{C,\epsilon}:(\overline{B}^{k}V_{\alpha},\widehat{\ell}_{\epsilon})\to{\mathbb{Q}}$ defined using $p^{k,0}_{C,\epsilon}$. When $X$ is an exact cobordism from $Y_{-}$ to $Y_{+}$, let $C_{-},C_{+}$ be two closed singular chains in $Y_{-},Y_{+}$, which define two pointed morphism $p_{C_{-}},p_{C_{+}}$. If $C_{-}$ and $C_{+}$ are homologous in $X$, then one can show that $p_{C_{-}},p_{C_{+}}$ and the $BL_{\infty}$ morphisms induced from $X$ are compatible in the sense of Definition 2.11 by counting holomorphic curves in $\widehat{X}$ with a point passing through the singular chain whose boundary is $-C_{-}\sqcup C_{+}$. ^$\widehat{\ell}_{C,\epsilon}$ up to homotopy only depends on the homology class $[C]$. Therefore for any $k\geq 1$, we have a linear map
$\delta^{\vee}_{\epsilon}:H_{*}(\overline{B}^{k}V_{\alpha},\widehat{\ell}_{\epsilon})\otimes H_{*}(Y,{\mathbb{Q}})\to{\mathbb{Q}},\quad(x,[C])\mapsto\widehat{\ell}_{C,\epsilon}(x).$
Or equivalently, we can write it as $\delta_{\epsilon}:H_{*}(\overline{B}^{k}V_{\alpha},\widehat{\ell}_{\epsilon})\to H^{*}(Y,{\mathbb{Q}})$. Combining the method in [6] and the argument in §5 in [31], one can show that $H_{*}(\overline{B}^{1}V_{\alpha},\widehat{\ell}_{\epsilon})\to H^{*}(Y,{\mathbb{Q}})$ is isomorphic to the composition of the maps $SH_{+,S^{1}}^{*}(W;{\mathbb{Q}})\to H^{*}_{S^{1}}(W;{\mathbb{Q}}):=H^{*+1}(W;{\mathbb{Q}})\otimes_{\mathbb{Q}}{\mathbb{Q}}[U,U^{-1}]/[U]\to H^{*+1}(Y;{\mathbb{Q}})$, where $W$ is an exact filling and $\epsilon$ is the augmentation induced from the filling $W$ [^5] Note that symplectic cohomology in [38] is graded by $n-\mu_{CZ}$, this explains the discrepancy of parity.. This fact was used in [37, 38] (but not phrased in SFT) to define obstructions to Weinstein fillings. In principle, $H_{*}(\overline{B}^{k}V_{\alpha},\widehat{\ell}_{\epsilon})\to H^{*}(Y,{\mathbb{Q}})$ can be used to obstruct Weinstein fillings if the image contains an element of degree $>\frac{\dim Y+1}{2}$ for all possible augmentations that could be from a Weinstein filling, e.g. ${\mathbb{Z}}$-graded augmentations if $c_{1}(Y)=0$ and $\dim Y\geq 5$.
In the following, we will explain the functorial part of $\delta_{\epsilon}$ without proof. Given an exact cobordism $X$ from $Y_{-}$ to $Y_{+}$, let $C$ be a closed chain in $X$. Then by counting rational holomorphic curves in $X$ with a marked point mapped to $C$, we obtain a family of maps $\phi_{C}^{k,l}:S^{k}V_{\alpha_{+}}\to S^{l}V_{\alpha_{-}}$. Along with the $BL_{\infty}$ morphism $\phi^{k,l}$ from $X$, we can construct a map $\widehat{\phi}_{C}:EV_{\alpha_{+}}\to EV_{\alpha_{-}}$ from $\phi^{k,l}_{C},\phi^{k,l}$ by the same rule of $\widehat{\phi}_{\bullet}$. Now that $C$ is closed, we have $\widehat{\phi}_{C}\circ\widehat{p}_{+}=\widehat{p}_{-}\circ\widehat{\phi}_{C}$. Then given an augmentation $\epsilon$ of $V_{\alpha_{-}}$, we have a linearized relation $\widehat{\phi}_{C,\epsilon}\circ\widehat{p}_{+,\epsilon\circ\phi}=\widehat{p}_{-,\epsilon}\circ\widehat{\phi}_{C,\epsilon}$, where $\widehat{\phi}_{C,\epsilon}=\widehat{F}_{\epsilon}\circ\widehat{\phi}_{C}\circ\widehat{F}_{-\epsilon\circ\phi}$. In particular, $\widehat{\phi}^{k,0}_{C,\epsilon}$ defines a chain map $(\overline{B}^{k}V_{\alpha_{+}},\widehat{\ell}_{\epsilon\circ\phi})\to{\mathbb{Q}}$. By a similar argument as before, such construction yields a map
$\delta^{\vee}_{X,\epsilon}:H_{*}(\overline{B}^{k}V_{\alpha_{+}},\widehat{\ell}_{\epsilon\circ\phi})\otimes H_{*}(X;{\mathbb{Q}})\to{\mathbb{Q}}$
where the dual version $H_{*}(\overline{B}^{k}V_{\alpha_{+}},\widehat{\ell}_{\epsilon\circ\phi})\to H^{*}(X;{\mathbb{Q}})$ is denoted by $\delta_{X,\epsilon}$. If $C$ comes from $H_{*}(Y_{+})$, then $\delta^{\vee}_{X,\epsilon}(-\otimes C)$ coincides with $\delta^{\vee}_{\epsilon\circ\phi}(-\otimes C)$ by the same argument of Proposition 5.14 in [31]. If $C$ comes from $Y_{-}$, then $\delta^{\vee}_{X,\epsilon}$ factors through $\widehat{\phi}_{\epsilon}:H_{*}(\overline{B}^{k}V_{\alpha_{+}},\widehat{\ell}_{\epsilon\circ\phi})\to H_{*}(\overline{B}^{k}V_{\alpha_{-}},\widehat{\ell}_{\epsilon})$ and $\delta^{\vee}_{X,\epsilon}(-\otimes C)=\delta^{\vee}_{\epsilon}(\widehat{\phi}_{\epsilon}(-)\otimes C)$ by an argument similar to Proposition 5.14 in [31]. Dualizing those properties, we have the following commutative diagram,
[…]
5.1.2. Tangency conditions
Another type of generalization is considering point constraints with tangency conditions, i.e. we consider curves in the symplectization passing through a fixed point $p$ and tangent to a local divisor near $p$ with order $m$. Such holomorphic curves were considered in [9, 34]. Those curves also give rise to pointed morphisms and hence can be used to define a new order. In many cases, if we have a holomorphic curve with a point constraint without tangent conditions, then multiple covers of it might have tangent properties. One can show that the order with a tangency condition for $(S^{2n-1},\xi_{std}),n\geq 2$ is always $1$ no matter what is the order of tangency. The curve is easy to find in a very thin ellipsoid with one positive puncture asymptotic to a multiple cover of the shortest Reeb orbit. Then one can use a neck-stretching argument to prove that it is independent of the augmentation. The invariants with local tangent constraints for exact domains are defined and computed in [19].
5.1.3. Multiple point constraints
It is natural to consider generalizations of pointed morphisms of $BL_{\infty}$ algebras to maps induced from counting curves with multiple constraints. For example, we can consider rational holomorphic curves with $2$ marked points passing through two fixed points in the contact manifold, where the curve can be disconnected. More specifically, we have three families of maps from $S^{k}V$ to $S^{l}V$ that are $p_{\bullet\bullet}^{k,l}$ coming from counting connected holomorphic curves with two marked points, $p_{\bullet_{1}}^{k,l},p_{\bullet_{2}}^{k,l}$ coming from counting connected holomorphic curves with each one of the point constraints respectively. Then we can assemble them to $\widehat{p}_{\bullet\bullet}$ by the same rule of $\widehat{p}_{\bullet}$ except the middle level consists of one $p_{\bullet\bullet}$ or both $p_{\bullet_{1}},p_{\bullet_{2}}$. Note that the combinatorics behind $\widehat{p}_{\bullet\bullet}$ is slightly different from $\widehat{p}_{\bullet}$ and $\widehat{p}$. Nevertheless, we have $\widehat{p}_{\bullet\bullet}\circ\widehat{p}=\widehat{p}\circ\widehat{p}_{\bullet\bullet}$ and given a $BL_{\infty}$ augmentation, we can similarly define $\widehat{p}_{\bullet\bullet,\epsilon}$, which is assembled from $p_{\bullet\bullet,\epsilon},p_{\bullet_{1},\epsilon}$ and $p_{\bullet_{2},\epsilon}$. In general, we can consider disconnected rational holomorphic curves with $k$ marked points passing through $k$ point constraints, which give rise to an operator $\widehat{p}_{k\bullet}$ on $EV$. $\widehat{p}_{k\bullet}$ is homotopic to $\widehat{p}_{\bullet}^{k}$.
For a pointed morphism, we can define a slightly different order using the full linearized $BL_{\infty}$ structure instead of just the $L_{\infty}$ structure as follows.
Definition 5.1.
Given a $BL_{\infty}$ augmentation and a pointed map $p_{\bullet}$, we define
$\tilde{O}(V,\epsilon,p_{\bullet}):=\min\left\{k\left|1\in\pi_{{\mathbb{Q}}}\circ\widehat{p}_{\bullet,\epsilon}|_{H_{*}(\overline{B}^{k}\overline{S}V,\widehat{p}_{\epsilon})}\right.\right\},$
where $\pi_{{\mathbb{Q}}}:EV\to{\mathbb{Q}}\subset SV$ is the projection.
Proposition 5.2.
$O(V,\epsilon,p_{\bullet})\leq\tilde{O}(V,\epsilon,p_{\bullet})$.
Proof.
We use $\overline{B}^{k}_{2}\subset\overline{B}^{k}\overline{S}V$ to denote the subspace $(\bigoplus_{i=2}^{\infty}S^{i}V)\odot^{k-1}\overline{S}V$. Since $p^{k,0}_{\epsilon}=0$, we have $\overline{B}^{k}_{2}$ is a subcomplex. Moreover, the quotient complex $\overline{B}^{k}\overline{S}V/\overline{B}^{k}_{2}$ is exactly $(\overline{B}^{k}V,\widehat{\ell}_{\epsilon})$. Note that $\pi_{{\mathbb{Q}}}\circ\widehat{p}_{\bullet,\epsilon}$ on $\overline{B}^{k}_{2}$ is zero, as $\widehat{p}_{\bullet,\epsilon}|_{\overline{B}^{k}_{2}}$ contains at least one $V$ component in the output. Moreover, $\pi_{{\mathbb{Q}}}\circ\widehat{p}_{\bullet,\epsilon}$ on $\overline{B}^{k}\overline{S}V/\overline{B}^{k}_{2}V$ is identified with ^$\widehat{\ell}_{\bullet,\epsilon}$ on $\overline{B}^{k}V$. Therefore $1\in\pi_{{\mathbb{Q}}}\circ\widehat{p}_{\bullet,\epsilon}|_{H^{*}(\overline{B}^{k}\overline{S}V,\widehat{p}_{\epsilon})}$ implies that $1\in\operatorname{im}\widehat{\ell}_{\bullet,\epsilon}|_{H_{*}(\overline{B}^{k}V,\widehat{\ell}_{\epsilon})}$. Hence the claim follows. ∎
Similarly, we can define
$\tilde{O}(V,\epsilon,p_{k\bullet}):=\min\left\{k\left|1\in\pi_{{\mathbb{Q}}}\circ\widehat{p}_{k\bullet,\epsilon}|_{H_{*}(\overline{B}^{k}\overline{S}V,\widehat{p}_{\epsilon})}\right.\right\}.$
To give the analog of $O(V_{\alpha},\epsilon,p_{\bullet})$ when we have multiple point constraints, we consider the projection $\pi_{m}:\overline{B}^{k}\overline{S}V_{\alpha}\to\overline{B}^{k}\overline{B}^{m}V_{\alpha}$. Then we have a filtration $\overline{B}^{k}\overline{S}V_{\alpha}:=\ker\pi_{0}\supset\ker\pi_{1}\supset\ldots$. Then we define the width of an element in $\overline{B}^{k}\overline{S}V_{\alpha}$ by
$w(v)=\sup\left\{m|v\in\ker\pi_{m}\right\}.$
We claim that $w(\widehat{p}_{\epsilon}(v))\geq w(v)$. To see this, we will fist prove $\pi_{m}\circ\widehat{p}_{\epsilon}=\pi_{m}\circ\widehat{p}_{\epsilon}\circ\pi_{m}$ as follows. For $x\in\overline{B}^{k}\overline{S}V_{\alpha}$, we can write $x=x_{0}+x_{1}$ for $x_{0}=\pi_{m}(x)\in\overline{B}^{k}\overline{B}^{m}V_{\alpha},x_{1}\in\ker\pi_{m}$. Since $p^{k,0}_{\epsilon}=0$, we have $\pi_{m}\circ\widehat{p}_{\epsilon}(x_{1})=0$. As a consequence, we have $\pi_{m}\circ\widehat{p}_{\epsilon}(x)=\pi_{m}\circ\widehat{p}_{\epsilon}\circ\pi_{m}(x)$. Therefore if $\pi_{m}(v)=0$, we have $\pi_{m}\circ\widehat{p}_{\epsilon}(v)=0$. In particular, $w(\widehat{p}_{\epsilon}(v))\geq w(v)$. Therefore $\pi_{m}\circ\widehat{p}_{\epsilon}|_{\overline{B}^{k}\overline{B}^{m}V_{\alpha}}$ squares to zero, hence is a differential on $\overline{B}^{k}\overline{B}^{m}V_{\alpha}$. $\pi_{m}\circ\widehat{p}_{\epsilon}$ uses the knowledge of $p^{k,l}_{\epsilon}$ for $l\leq m$ and $\pi_{1}\circ\widehat{p}_{\epsilon}=\widehat{\ell}_{\epsilon}$. On the other hand, given the operator $\widehat{p}_{m\bullet}$ defined using $m$ point constraints, we have the deformed chain map $\widehat{p}_{m\bullet,\epsilon}=\widehat{F}_{\epsilon}\circ\widehat{p}_{m\bullet}\circ\widehat{F}_{-\epsilon}$ from $(EV,\widehat{p}_{\epsilon})$ to itself. Now we consider $\widehat{p}_{m\bullet,\epsilon}$ on the subcomplex $(\overline{E}V:=\overline{S}\overline{S}V,\widehat{p}_{\epsilon})$. Since $\widehat{p}_{m\bullet,\epsilon}$ can have at most $m$ nontrivial connected components, we have $\pi_{{\mathbb{Q}}}\circ\widehat{p}_{m\bullet,\epsilon}(v)=0$ if $w(v)\geq m+1$. Hence we have $\pi_{{\mathbb{Q}}}\circ\widehat{p}_{m\bullet,\epsilon}\circ\pi_{m}\circ\widehat{p}_{\epsilon}=\pi_{{\mathbb{Q}}}\circ\widehat{p}_{m\bullet,\epsilon}\circ\widehat{p}_{\epsilon}=0$ and $\pi_{{\mathbb{Q}}}\circ\widehat{p}_{m\bullet,\epsilon}\circ\pi_{m}=\pi_{{\mathbb{Q}}}\circ\widehat{p}_{m\bullet,\epsilon}$, i.e. we have the following commutative diagram of chain complexes,
[…]
Definition 5.3.
We define $O(V,\epsilon,p_{m\bullet}):=\min\left\{k\left|1\in\operatorname{im}\pi_{{\mathbb{Q}}}\circ\widehat{p}_{m\bullet,\epsilon}|_{H_{*}(\overline{B}^{k}\overline{B}^{m}V,\pi_{m}\circ\widehat{p}_{\epsilon})}\right.\right\}$.
In particular, $O(V,\epsilon,p_{1\bullet})$ is Definition 2.10. Moreover, (5.1) implies the following.
Proposition 5.4.
$O(V,\epsilon,p_{m\bullet})\leq\tilde{O}(V,\epsilon,p_{m\bullet})$.
Given a strict exact filling $(W,\lambda)$ of $(Y,\alpha)$, assume $\widetilde{O}(V_{\alpha},\epsilon,p_{k\bullet})<\infty$ for the $BL_{\infty}$ augmentation $\epsilon$ coming from $W$. Then we can define the spectral invariant for $l\geq\widetilde{O}(V_{\alpha},\epsilon,p_{k\bullet})$,
$\mathfrak{r}^{\leq l}\langle\underbrace{p,\ldots,p}_{k}\rangle:=\inf\left\{a:T^{a}\in\operatorname{im}\pi_{\Lambda_{0}}\circ\widehat{p}_{k\bullet,\epsilon}|_{H_{*}(\overline{B}^{l}\overline{S}V^{\geq 0}_{\alpha},\widehat{p}_{\epsilon})}\right\}<\infty,$
where $V^{\geq 0}_{\alpha}$ is the free $\Lambda_{0}:=\{\sum a_{i}T^{b_{i}}|b_{i}\geq 0,\lim b_{i}=\infty,a_{i}\in{\mathbb{Q}}\}$ module generated by good Reeb orbits of $\alpha$. This is exactly the higher symplectic capacity $\mathfrak{r}^{\leq l}\langle\underbrace{p,\ldots,p}_{k}\rangle$ defined by Siegel in [34, §6]. In [34], the author considered disconnected holomorphic curves in $W$ with $k$ point constraints and viewed them as a chain map from $\overline{S}\operatorname{CHA}(Y)=EV_{\alpha}$ to $\Lambda$. Then the capacity is defined to be the infimum of $a$ such that there is a closed class in $x\in\overline{B}^{k}SV^{\geq 0}_{\alpha}$ such that $x$ is mapped to $T^{a}$ by the chain map. The equivalence of these two definitions can be seen from the same argument of Proposition 5.14 in [31]. Similarly, combining the tangency conditions and multiple point constraints, we can define the analogous orders, whose spectral invariant is again equivalent to the higher capacity $\mathfrak{r}^{\leq l}\langle{\mathcal{T}}^{m_{1}}p,\ldots,{\mathcal{T}}^{m_{k}}p\rangle$ in [34, §6]. Similarly, the spectral invariant for $O(V,\epsilon,p_{\bullet})$ is equivalent to $\mathfrak{g}^{\leq l}_{1}$ and the spectral invariant for the analogous version of pointed morphism with tangency conditions of order $m$ is $\mathfrak{g}^{\leq l}_{t^{m}}$ in [34].
Example 5.5.
Let $W$ be an irrational ellipsoid with $\alpha$ the contact form on $\partial W$. Then we have $\tilde{O}(V_{\alpha},\epsilon_{W},p_{k\bullet})=1$ for all $k\geq 1$. To see this, we let $\gamma$ denote the shortest Reeb orbits, then for a generic point in $W$ and a generic admissible almost complex structure, there is one [^6] More precisely, the algebraic count of such curves is $1$. holomorphic plane in $\widehat{W}$ asymptotic to $\gamma$ and passing through the point. Since $\mu_{CZ}(\gamma^{\prime})+n-3=0\mod 2$ for any Reeb orbit $\gamma^{\prime}$, we have $q_{\gamma}\in\overline{B}^{1}\overline{S}V_{\alpha}$ is a closed class and $\pi_{\mathbb{Q}}\circ\widehat{p}_{\bullet,\epsilon_{W}}(q_{\gamma})=1$ by the holomorphic curve above. In general, we have $q^{k}_{\gamma}\in\overline{B}^{1}\overline{S}V_{\alpha}$ is closed by degree reasons. On the other hand, $\pi_{\mathbb{Q}}\circ\widehat{p}_{k\bullet,\epsilon_{W}}|_{\overline{S}^{k}V_{\alpha}\subset\overline{B}^{1}\overline{S}V_{\alpha}}$ must only use disconnected curves with $k$ components and each component has one positive puncture, for otherwise, genus has to be created. As a consequence, we have $\pi_{\mathbb{Q}}\circ\widehat{p}_{k\bullet,\epsilon_{W}}(q^{k}_{\gamma})=1$ by the disconnected $k$ copies of the curve above. Then one can show such curve is independent of the augmentation, as the curve above is contained in the symplectization after a neck-stretching, see [36, 37]. Although there is nothing interesting happening for the order $\tilde{O}(V_{\alpha},\epsilon_{W},p_{k\bullet})$. We know that the spectral invariant $\mathfrak{r}^{\leq l}\langle\underbrace{p,\ldots,p}_{k}\rangle$ is defined for all $k\geq 1$ and $l\geq 1=\tilde{O}(V_{\alpha},\epsilon_{W},p_{k\bullet})$. Those numerical invariants are very sensitive to the shape of $W$ and are powerful tools to study embedding problems, see [34] for details.
Remark 5.6.
The spectral invariants for $O(V,\epsilon,p_{m\bullet})$ provide a new family of higher symplectic capacities.
5.2. Higher genera
Another natural direction to generalize is by increasing the genus of holomorphic curves, i.e. considering the full SFT. Originally, the full SFT was phrased as a differential Weyl algebra with a distinguished odd degree Hamiltonian ${\bm{H}}$ such that ${\bm{H}}\star{\bm{H}}=0$ [13]. There are two closely related ways to view the full SFT as a functor from ${\mathfrak{Con}}$ in a more convenient way, namely the $BV_{\infty}$ formulation [11] and the $IBL_{\infty}$ formulation [10]. In the following, we will first briefly recall their definitions. In view of the $BL_{\infty}$ formalism in this paper, we will give a slightly different but equivalent definition of $IBL_{\infty}$ algebras. Then we will make some speculations assuming the analytical foundation of the full SFT is completed.
5.2.1. $IBL_{\infty}$ algebras and $BV_{\infty}$ algebras
The original definition of $IBL_{\infty}$ algebra on a ${\mathbb{Z}}/2$ graded vector space involves taking a suspension as the formalism for $L_{\infty}$ algebras in [31]. To align with the notation of $BL_{\infty}$ algebras, we will not take the suspension. In particular, if $V$ is an $IBL_{\infty}$ algebra in Definition 5.7, then $V[-1]$ is an $IBL_{\infty}$ algebra in the sense of [10]. Let $\phi:S^{k}V\to S^{l}V$, we can define $\widehat{\phi}:\overline{S}V\to\overline{S}V$ by $\widehat{\phi}=0$ on $S^{m}V$ with $m<k$ and
$\widehat{\phi}(v_{1}\ldots v_{m})=\sum_{\sigma\in Sh(k,m-k)}\frac{(-1)^{\diamond}}{k!(m-k)!}\phi(v_{\sigma(1)}\ldots v_{\sigma(k)})v_{\sigma(k+1)}\ldots v_{\sigma(m)},$
for $m>k$.
Definition 5.7 ([10, Definition 2.3]).
Let $V$ be a ${\mathbb{Z}}/2$ graded vector space over ${\bm{k}}$, an $IBL_{\infty}$ structure on $V$ is a family of operators $p^{k,l,g}:S^{k}V\to S^{l}V$ for $k,l\geq 1$ and $g\geq 0$, such that
$\widehat{p}:=\sum_{k,l=1}^{\infty}\sum_{g=0}^{\infty}\widehat{p}^{k,l,g}\hbar^{k+g-1}\tau^{k+l+2g-2}:\overline{S}V[[\hbar,\tau]]\to\overline{S}V[[\hbar,\tau]]$
satisfies that $\widehat{p}\circ\widehat{p}=0$ and $|\widehat{p}|=1$. Here $|\hbar|=0$ and $|\tau|=0$.
In the case of $V$ being ${\mathbb{Z}}$ graded, then one can define $IBL_{\infty}$ structures of degree $d$ by requiring $|\hbar|=2d$ and $|\widehat{p}|=-1$. In the case of SFT, $d=n-3$ for a $12n-$ 1 dimensional contact manifold $(Y,\xi)$ with $c_{1}(\xi)=0$. If moreover we know that for any $v_{1},\ldots,v_{k}\in V$ and $g\geq 0$ there exist at most finitely many $l\geq 1$, such that $p^{k,l,g}(v_{1}\ldots v_{k})\neq 0$, then $\widehat{p}$ is well-defined on $\overline{S}V[[h]]$ by setting $\tau=1$. Note that we only consider $p^{k,l,g}$ with $l\geq 1$ in Definition 5.7. In the context of SFT, this is equivalent to that there is no holomorphic curve with no negative punctures in the symplectization, which is in general not true. Indeed, Definition 5.7 is used to model the linearized theory, i.e. SFT associated to an exact domain. In general, we will need the following (cocurved) $IBL_{\infty}$ algebras.
Definition 5.8.
Let $V$ be a ${\mathbb{Z}}_{2}$ graded vector space over ${\bm{k}}$, a (cocurved) $IBL_{\infty}$ structure on $V$ is a family of operators $p^{k,l,g}:S^{k}V\to S^{l}V$ for $k\geq 1,l,g\geq 0$, such that
  1. (1)
    for $v_{1},\ldots,v_{k}\in V,g\geq 0$, there are finitely many $l$ such that $p^{k,l,g}(v_{1}\ldots v_{k})\neq 0$,
  2. (2)
    $\widehat{p}:=\sum_{k=1}^{\infty}\sum_{l,g=0}^{\infty}\widehat{p}^{k,l,g}\hbar^{k+g-1}:SV[[\hbar]]\to SV[[\hbar]]$ satisfies $\widehat{p}\circ\widehat{p}=0$ and $|\widehat{p}|=1$, where $\widehat{p}(1)$ is defined to be 00.
Note that Definition 5.8 gives rise to a $BV_{\infty}$ algebra introduced in [11], which was used to define algebraic torsions in [25]. For the translations between $BV_{\infty}$ algebras and differential Weyl algebras with distinguished Hamiltonians, see [11] for details. In the following, we give an alternative description of Definition 5.8. Assume we are given a family of operators operators $p^{k,l,g}:S^{k}V\to S^{l}V$ for $k\geq 1,l,g\geq 0$, such that (1) of Definition 5.8 is satisfied. We use $EV[[h]]$ to denote $\overline{S}(SV[[h]])$, where $\overline{S}$ is the symmetric product over ${\bm{k}}[[\hbar]]$. Then we define $\widehat{p}:EV[[h]]\to EV[[h]]$ by the following graph description. We use a tree with $k+1$ vertices with a label $g\geq 0$ to represent a class $\hbar^{g}v_{1}\ldots v_{k}\in SV[[\hbar]]$. Then we use the graph with to represent operators $p^{k,l,g}$, but we label the with a number $g\geq 0$. To define $\widehat{p}$, we represent a class in $EV[[\hbar]]$ by a row of trees with labels, then we glue in one graph representing $p^{k,l,g}$ and dashed vertical lines representing the identity map. Here we allow cycles to be created. The rule for the output is the same as before with the degree of $\hbar$ determined by the sum of the labels $g$ in the connected component of the glued graph with the number of (independent) cycles in that component. Then $\widehat{p}$ is the sum of all such glued graphs.
$g_{1}$$g_{2}$$g_{3}$$g$
Figure 8. A component of $\widehat{p}$ from $\hbar^{g_{1}}S^{3}V\odot\hbar^{g_{2}}S^{3}V\odot\hbar^{g_{3}}S^{2}V$ to $\hbar^{g_{1}+g_{2}+g+1}S^{6}V\odot\hbar^{g_{3}}S^{2}V$ using $p^{3,3,g}$.
One can think of the tree representing a class in $SV[[\hbar]]$ as a genus $g$ surface with $k$ negative punctures and $p^{k,l,g}$ represented by a genus $g$ surface with $k$ positive punctures and $l$ negative punctures. Then a glued graph represented a possibly disconnected surface with only negative punctures, which represents a class in $EV[[\hbar]]$ with the degree of $\hbar$ is the sum of the genus of all connected components.
Definition 5.9.
$(V,p^{k,l,g})$ is a (curved) $IBL_{\infty}$ algebra iff $\widehat{p}:EV[[h]]\to EV[[h]]$ defined above satisfies $\widehat{p}\circ\widehat{p}=0$ and $|\widehat{p}|=1$.
Proposition 5.10.
Definitions 5.8 and 5.9 are equivalent.
Proof.
Note that $\widehat{p}:SV[[\hbar]]\subset EV[[\hbar]]\to SV[[\hbar]]$ as we can not increase the number of connected components in the graph description of $\widehat{p}$. However $\widehat{p}|_{SV[[\hbar]]}$ is exactly the $\widehat{p}$ in Definition 5.8 as the $\hbar^{k-1}$ is exactly the number of cycles in the glued graph. Therefore Definition 5.9 implies Definition 5.8. On the other hand, assume $\widehat{p}:SV[[\hbar]]\to SV[[\hbar]]$ squares to zero. Consider the $\hbar^{N}S^{m}V$ part of $\widehat{p}\circ\widehat{p}(v_{1}\ldots v_{n})$ for $v_{i}\in V$, which is zero. Since $\widehat{p}$ on $SV[[\hbar]]$ has the graph description on one tree as explained above, we need to consider possibly disconnected graphs with $n$ input vertices and $m$ output vertices and two vertices such that if we glue all input vertices together by adding a new vertex, the resulted graph has $N$ cycles. Then there are two cases, (1) the two vertices are in different components, then all of such graphs will pair up and cancel each other by $|\widehat{p}|=1$; (2) the two vertices are in the same component with the remaining components are compositions of dashed lines. Assume the component with has $a$ inputs and $b$ outputs and has genus $k$. We will call $(a,b,k)$ the signature of the glued graph. Then we must have $n-a=m-b$ is the number of components from dashed lines and $a+k-1=N$. We use $p_{2}^{a,b,k}$ to denote the map from $S^{a}V$ to $S^{b}V$ defined by two components such that the glued graph with signature $(a,b,k)$. Or equivalently, $p_{2}^{a,b,k}$ is defined by all possible two-level breakings of a graph with signature $(a,b,k)$. We formally define $p^{a,b,k}=0$ if $b<0$, then the vanishing of the $\hbar^{N}S^{m}V$ part of $\widehat{p}\circ\widehat{p}(v_{1}\ldots v_{n})$ implies
$\sum_{a=1}^{n}p_{2}^{a,m-n+a,N-a+1}\ast^{n-a}\operatorname{id}=0,$
for all $n\geq 1,m,N\geq 0$. By setting $n=1$, we have $p_{2}^{1,m,N}=0$ for $m,N\geq 0$. Then by setting $n=2$ and $p_{2}^{1,m,N}=0$, we have $p_{2}^{2,m,N}=0$. Similarly we have $p_{2}^{n,m,N}=0$ for all $n\geq 1,m,N\geq 0$. This is exactly describing all maps from two-level breakings of a connected graph with signature $(n,m,N)$ should sum up to zero. It implies that $\widehat{p}^{2}=0$ on $EV[[\hbar]]$ by the same argument in Proposition 2.8 in [31]. ∎
Remark 5.11.
From the proof of Proposition 5.10, both Definitions 5.8 and 5.9 are equivalent to $p_{2}^{n,m,N}=0$, which in the SFT world, corresponds to the algebraic counts of all rigid codimension $1$ breaking of connected holomorphic curves with $n$ positive punctures and $m$ negative punctures and genus $N$ sum up to zero.
Note that $SV\subset SV[[h]]$ induces an inclusion $EV\subset EV[[h]]$, and we use $\pi_{0}$ to denote the natural projection $EV[[\hbar]]\to EV$. It is easy to check that if $x\in\ker\pi_{0}$, then $\widehat{p}(x)\in\ker\pi_{0}$. As a consequence, $\pi_{0}\circ\widehat{p}|_{EV}:EV\to EV$ squares to zero. Moreover, $\pi_{0}\circ\widehat{p}|_{EV}$ is assembled from $p^{k,l,0}$. Then we have the following instant corollary.
Corollary 5.12.
Let $(V,p^{k,l,g})$ be a (cocurved) $IBL_{\infty}$ algebra, then $(V,p^{k,l}:=p^{k,l,0})$ is a $BL_{\infty}$ algebra.
5.2.2. A grid of torsions
Given a family of operators $\phi^{k,l,g}:S^{k}V\to S^{l}V^{\prime}$ for $k\geq 1,l,g\geq 0$, assume that for any $g\geq 0$ and $v_{1},\ldots,v_{k}\in V$, there are at most finitely many $l$ such that $\phi^{k,l,g}(v_{1}\ldots v_{k})\neq 0$. Then we can assemble $\widehat{\phi}:EV[[\hbar]]\to EV^{\prime}[[\hbar]]$ by the same rule for $BL_{\infty}$ morphisms except that cycles are allowed to be created and the rule for the order of $\hbar$ is the same as for $\widehat{p}$.
Definition 5.13.
The family of operators is an $IBL_{\infty}$ morphism from $(V,p^{k,l,g})$ to $(V^{\prime},q^{k,l,g})$ iff $\widehat{q}\circ\widehat{\phi}=\widehat{\phi}\circ\widehat{p}$ and $|\widehat{\phi}|=0$.
Then we have a trivial $IBL_{\infty}$ algebra $\mathbf{0}:=\{0\}$ with $p^{k,l,g}=0$. $\mathbf{0}$ is an initial object in the category of $IBL_{\infty}$ algebras with $\phi^{k,l,g}=0$. Then an $IBL_{\infty}$ augmentation of $V$ is an $IBL_{\infty}$ morphism from $V$ to $\mathbf{0}$. $IBL_{\infty}$ augmentations may not always exist. One obstruction is torsion. Unlike the torsions for $BL_{\infty}$ algebras, there are many more torsions for $IBL_{\infty}$ algebras. Let $E^{k}V[[\hbar]]:=\overline{B}^{k}SV[[\hbar]]$.
Definition 5.14.
For $n,m\geq 0$, we say $V$ has a $(n,m)$ torsion if $[\hbar^{n}]=0\in H_{*}(E^{m+1}V[[\hbar]])$.
Then the algebraic torsion in [25] is the $(n,0)$ torsion of the $IBL_{\infty}$ algebra associated to a contact manifold by SFT.
Proposition 5.15.
If $V$ has $(n,m)$ torsion, then $V$ has $(n+m,0)$ torsion.
Proof.
We use $\hbar^{-k}SV[[\hbar]]$ to denote $SV[[\hbar]]\otimes_{{\bm{k}}[[\hbar]]}\hbar^{-k}{\bm{k}}[[\hbar]]$, where $\hbar^{-k}{\bm{k}}[[\hbar]]$ is the ${\bm{k}}[[\hbar]]$ module generated by $\hbar^{-k}$. Then we have a map $c_{k}:S^{k}SV[[\hbar]]\to\hbar^{-k+1}SV[[\hbar]]$ for $k\geq 2$ defined by
$w_{1}\odot\ldots\odot w_{k}\mapsto\hbar^{-k+1}w_{1}\ldots w_{k}.$
Pictorially, $c_{k}$ is simply connecting all trees compensated with $\hbar^{-k+1}$. Let $\iota$ be the inclusion $SV[[\hbar]]\to h^{-k+1}SV[[\hbar]]$. We claim
$C_{m}:=\iota+\sum_{i=2}^{m}c_{i}$
defines a chain map from $E^{m}V[[\hbar]]$ to $\hbar^{-m+1}SV[[\hbar]]$ for $m\geq 2$. Assume the claim was proven, if $V$ admits a $(n,m)$ torsion, i.e. $\widehat{p}(x)=\hbar^{n}$ for some $x\in E^{m+1}V[[\hbar]]$. Since $C_{m+1}|_{SV[[\hbar]]}=\iota$, in particular, we have $C_{m+1}(\hbar^{n})=\hbar^{n}$. Hence we know that $\widehat{p}\circ C_{m+1}(x)=\hbar^{n}$. Then $\widehat{p}(\hbar^{m}C_{m+1}(x))=\hbar^{n+m}$ and $\hbar^{m}C_{m+1}(x)\in SV[[\hbar]]$, i.e. $V$ carries a $(n+m,0)$-torsion. To prove the claim, we will show that $C_{m}\circ\widehat{p}=\widehat{p}\circ C_{m}$ on $S^{i}SV[[\hbar]]$ for each $1\leq i\leq m$. Using the graph description, $\widehat{p}$ on $S^{i}SV[[\hbar]]$ is decomposed w.r.t. the subset $I$ of $\{1,\ldots,i\}$ such that the nontrivial component of $\widehat{p}$ is glued to clusters indexed by $I$. Let $\widehat{p}_{j}$ denote the component such that the subset has $j$ elements, then $\widehat{p}|_{S^{i}SV[[\hbar]]}=\sum_{j=1}^{i}\widehat{p}_{i}$. The graph description of $C_{m}\circ\widehat{p}_{j}$ is simply connecting all components of $\widehat{p}_{j}$ multiplied by $\hbar^{j-i}$. One the other hand, in the graph description of $\widehat{p}\circ C_{m}|_{S^{i}SV[[\hbar]]}$, we have the same graph. The order of $\hbar$ (in addition to the genus associated to the graph of $\widehat{p}_{j}$) is $\hbar^{-i+1}$ from $C_{m}$ on $S^{i}SV[[\hbar]]$ multiplied by $\hbar^{j-1}$ from the extra genus when we glue $\widehat{p}_{j}$ to the one combined tree compared to original $i$ trees. In particular, we have $C_{m}\circ\widehat{p}=\widehat{p}\circ C_{m}$ on $S^{i}SV[[\hbar]]$. Hence the claim follows. ∎
However, the torsion for the associated $BL_{\infty}$ algebra is not necessarily a $(0,m)$ torsion unless $m=0$ [5]. We use $p_{0}$ to denote the genus 00 part of $p$, i.e. the induced $BL_{\infty}$ algebra on $V$. Indeed if $1=\widehat{p_{0}}(x)$ for $x\in E^{m+1}V$, then we only have $\widehat{p}(x)=1+O(\hbar)$, where $O(\hbar)\in\ker\pi_{0}$ with $O(\hbar)$ potentially coming from holomorphic curves with higher genera. On the other hand, as we will see later, $(0,m)$ torsion always implies $m$ torsion for the $BL_{\infty}$ algebra. To explain the relations, we will explain the following procedure which increases genus one at a time.
Let $(V,p^{k,l,g})$ be an $IBL_{\infty}$ algebra. We define $EV[[\hbar]]_{m}:=EV[[\hbar]]\otimes_{{\bm{k}}[[\hbar]]}({\bm{k}}[[\hbar]]/[\hbar^{m+1}])$, then we consider the projection $\pi_{m}:EV[[\hbar]]\to EV[[\hbar]]_{m}$. Then we have a filtration
$EV[[\hbar]]:=\ker\pi_{-1}\supset\ker\pi_{0}\supset\ker\pi_{1}\supset\ldots$
We define the $\hbar$-width $w_{\hbar}(v)\in{\mathbb{N}}\cup\{\infty\}$ for $v\in EV[[\hbar]]$ to be
$\sup\left\{m|v\in\ker\pi_{m}\right\}+1.$
We can view ${\bm{k}}[[\hbar]]/[\hbar^{m+1}]$ as polynomials of $\hbar$ of degree at most $m$, in particular $EV[[\hbar]]_{m}$ can be viewed as a subspace of $EV[[\hbar]]$. Then for $x\in EV[[\hbar]]$, we have $x=\pi_{m}x+x_{1}$ for $x_{1}\in\ker\pi_{m}$. It is direct to check that $\pi_{m}\circ\widehat{p}(x_{1})=0$. As a consequence, we have $\pi_{m}\circ\widehat{p}\circ\pi_{m}=\pi_{m}\circ\widehat{p}$. Then we have $w_{\hbar}(\widehat{p}(v))\geq w_{\hbar}(v)$ for all $v\in EV[[\hbar]]$. As a consequence, we have the following commutative diagram,
[…]
And we have $\widehat{p}_{m}\circ\widehat{p}_{m}=0$. If we unwrap the definition, then we know that $\widehat{p}_{m}$ uses $p^{k,l,g}$ for $g\leq m$. Take $m=1$ as an example, $\widehat{p}_{1}$ uses both $p^{k,l,1},p^{k,l,0}$. However, to get an output in $EV[[h]]_{1}$, the $k$ inputs of $p^{k,l,1}$ must glue to $k$ trees while at most $2$ of the $k$ inputs of $p^{k,l,0}$ can glue to the same tree. We can similarly define $E^{k}V[[\hbar]]_{m}$ and we have analogous descriptions.
Definition 5.16.
We say $(V,p^{k,l,g})$ has a $(n,m)_{k}$ torsion iff $[\hbar^{n}]=0\in H^{*}(E^{m+1}V[[\hbar]]_{k})$.
Then by definition, we always have $(n,m)_{k}$ torsion for $n>k$. Moreover, $(0,m)_{0}$ torsion is the $m$-torsion for $BL_{\infty}$ algebras. The $(n,m)$ torsion can be viewed as $(n,m)_{\infty}$ torsion. We summarize the basic properties of those torsions in the following.
Proposition 5.17.
The torsions have the following properties.
  1. (1)
    If $V$ has $(n,m)_{k}$ torsion then $V$ has $(n,m)_{k-1}$, $(n+1,m)_{k}$ and $(n,m+1)_{k}$ torsions. In particular if $V$ has $(n,m)$ torsion, then $V$ has $(n,m)_{k}$ torsion for any $k\geq 0$.
  2. (2)
    If $V$ has $(n,m)_{k}$ torsion, then $V$ has $(n+m,0)_{k}$ torsion.
Proof.
For (1), that $V$ carries $(n,m)_{k-1}$ torsion follows from (5.2). The existence of the other two torsions follows from the definition. (2) follows from the $EV[[\hbar]]_{k}$ version of Proposition 5.15. ∎
As a corollary, the existence of $(0,k)$ torsion implies both $k$-algebraic torsion and $k$-algebraic planar torsion. Contact manifolds in [31, Theorem 3.16 and 3.17] actually have $(0,k)$ torsion by the same argument as in [25]. Therefore it implies both algebraic torsion and algebraic planar torsion. Roughly speaking, different torsions make different requirements on holomorphic curves. For $(k,0)$ torsion, we can have higher genus curves without negative puncture to contribute to the torsion, while all higher genus curves with the same positive asymptotics and non-empty negative punctures must sum up to zero. For $(0,k)$ torsion, we can only have rational curves without negative puncture to contribute to the torsion, while all higher genus curves with the same positive asymptotics and non-empty negative punctures must sum up to zero. For $(0,k)_{0}$ torsion, we can only have rational curves without negative punctures to contribute to the torsion, and all rational curves with the same positive asymptotics and non-empty negative punctures must sum up to zero.
Let $2^{{\mathbb{N}}^{3}}$ denote the category of subsets of ${\mathbb{N}}^{2}\times({\mathbb{N}}\cup\{\infty\})$, the arrow from $V$ to $W$ is an inclusion $W\subset V$. $2^{{\mathbb{N}}^{3}}$ is a monoidal category where the monoidal structure is given by taking the union. We use $\operatorname{SFT}(Y)$ to denote the full SFT as an $IBL_{\infty}$ algebra in the sense of Definition 5.9 for a contact manifold $Y$.
Theorem 5.18.
Let $Y$ be a contact manifold, we define $\mathrm{T}(V):=\{(m,n,k)|\operatorname{SFT}(Y)\text{ has }(m,n)_{k}\text{ torsion}\}$. Then $\mathrm{T}:{\mathfrak{Con}}\to 2^{{\mathbb{N}}^{3}}$ is a covariant monoidal functor.
Remark 5.19.
For the proof of Theorem 5.18, we only need to construct $\operatorname{SFT}(Y)$ to the same extent as [31, Theorem 3.10], which is expected from [17, 23]. It is also expected that one can generalize the construction in [33], in particular, proving the existence of implicit atlases with cell-like stratification. One of the main differences is that gluing parameters are subject to more relations if we allow cycles to be created, and we may not be able to find a “free basis” to build a topological manifold with boundary like $(G_{II})_{T/}$ in the proof of [31, Theorem 3.10].
Remark 5.20 (Analog of planarity).
Given an $IBL_{\infty}$ augmentation $\epsilon$, then we have similar constructions of linearization to construct another $IBL_{\infty}$ structure $p^{k,l,g}_{\epsilon}$ such that $p^{k,l,g}_{\epsilon}=0$ whenever $l=0$. As a consequence, we arrive at an $IBL_{\infty}$ structure in the sense of Definition 5.7. Then we can introduce the analog of pointed morphisms and the analog of orders in the context of $IBL_{\infty}$ algebras, which in the SFT case considers holomorphic curves passing through a fixed point in the symplectization. Moreover, we will have a grid of orders as in the case of torsions above. However, it is a much harder task to find examples with holomorphic curves with higher genera.